A Complete Bibliography of Applied Statistics: 2000–2009

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

14 March 2019 Version 4.10

Title word cross-reference

2 × 2 [385]. α [155]. D [313]. κ [2, 139]. M [382]. N [155]. p [385].

-coefficients [139]. -estimation [382]. -optimal [313]. -values [385].

10th [291].

3-butadiene [132].

8 [299].

absence [12, 187]. abundance [118]. abuse [8, 310, 346]. accelerated [129]. accident [186]. accounting [342, 209]. accumulation [219]. accuracy [204, 171]. achievement [142]. acid [34, 205]. acquired [128, 79, 375]. across [277]. activity [277]. acupuncture [233]. acute [13]. Adaptive [312, 76, 309, 303]. additive [295, 404, 206, 56, 360, 223]. Adjusted [116, 196, 247]. adjustment [15, 209]. admission [15]. after [50, 186]. age [60, 149, 89, 335, 247]. age-adjusted [247]. age-stratified [60]. agents [384]. aggregated [141]. aggregation [243, 268]. agreement [166]. agricultural [276]. air [353, 372, 207]. airborne [347]. airline [113]. algorithm [313, 409, 349]. Algorithms [329]. alignments [389]. allocation [190, 380, 329]. alpha [85]. alpha-tocopherol [85]. amino [205]. amino-acid [205]. Analyses [37, 11, 49, 338]. Analysing [126, 98, 274, 8, 239, 38, 79, 309]. Analysis [107, 280, 63, 325, 397, 135, 74, 182, 243, 161, 360, 365, 151, 316, 318, 117, 46, 403, 381, 404, 258, 18, 10, 364, 390, 195, 231, 260, 51, 143, 286, 165, 27, 20, 344, 328, 168, 220, 24, 376, 222, 122, 19, 375, 163, 238, 44, 90, 147, 150, 350, 50, 268, 48, 183, 224, 388, 216, 386, 162, 160, 55, 259, 244, 339, 68, 36, 62, 179]. analytic [99, 221]. animal [51, 54]. Anisotropic [92]. annotations [269]. annual [265]. Anticipated [131]. Anticipating [140]. antiretroviral [189]. Application [131, 87, 106, 137, 232, 295, 275, 4, 212, 266, 277, 292, 1, 59, 373, 279, 327, 169, 152, 13, 27, 276, 355, 39, 79, 374, 237, 367, 376, 85, 84, 74, 28, 375, 337, 239, 402, 180, 282, 281, 322, 341, 53, 321, 365, 247, 409, 33]. Applications [107, 258, 312, 390, 253, 29, 3, 88, 165, 122, 75]. Applied [104, 119, 202, 220, 90, 388]. Approach [138, 235, 166, 14, 387, 26, 290, 79, 84, 115, 173, 65, 291, 38, 120, 129, 303]. approaches [344]. approximations [390]. archaeological [30]. area [15, 305, 65]. areal [311]. areas [250, 226]. arising [311]. arm [121]. artery [388]. arthritis [365]. artificial [175]. asexual [380]. assays [197, 254]. Assess [103]. Assessing [204, 194, 261, 142]. Assessment [124, 273, 256, 326, 405, 368, 173]. associated [212, 353, 372]. association [88, 224, 338]. associations [408]. assumptions [325]. Asthma [136]. asymmetric [20]. asymmetry [258]. Athens [124]. Atlantic [95]. atmospheric [186, 55]. atomic [213]. Attribution [12]. auctions [294, 340]. auditing [239]. augmented [181]. authenticity [253]. Author [322, 324]. authors [42, 78, 111, 148, 192, 251, 293, 330, 369, 413]. authorship [351]. auto [352]. auto-regressive [352]. autocorrelated [282]. automated [334]. automating [386]. autoregressive [52, 237, 211, 264, 33]. auxiliary [67]. average [352, 65, 61]. avoidance [158]. bacterial [389]. Baden [295]. Baden-Württemberg [295]. barrier [47]. base [259]. base-line [259]. Based [133, 109, 234, 256, 56, 393, 290, 400, 173, 65, 274, 188, 406, 360, 307, 298]. basic [60]. Bayesian [316, 295, 117, 189, 30, 10, 86, 15, 228, 114, 51, 56, 95, 326, 393, 165, 344, 39, 149, 374, 361, 302, 363, 245, 368, 222, 40, 399, 385, 380, 132, 227, 198, 310, 346, 350, 372, 203, 38, 186, 356, 9, 48, 171, 207, 129, 162, 80, 247, 205, 309, 383, 411, 179, 301]. bees [257]. behaviour [124, 307]. benefits [248]. Best [264]. beta [85, 323, 324, 296]. beta-binomial [323, 324, 296]. beta-carotene [85]. bets [329]. between [166, 142, 88, 153, 28, 380, 224]. beyond [279]. Bézier [327]. Bias [24, 204, 332, 93]. biased [190]. Binary [33, 58, 46, 287, 278, 376, 304, 402, 161, 356]. binomial [232, 392, 323, 9, 324, 296]. bioassay [54]. biological [75]. biomarker [410]. biomass [22]. Biplots [101, 318]. birth [267]. Bivariate [134, 218, 96, 402]. bomb [213]. bone [333]. Boolean [382]. bootstrap [142, 360]. bootstrap-based [360]. Boston [305, 353]. Both [130]. boundaries [91]. Bounded [133]. bovine [63, 229]. Box [19]. brain [312, 153]. Breast [104, 168, 90, 50, 179]. breathing [353]. British [8]. Building [81, 30]. butadiene [132].

Calculating [139, 349]. calculations [407]. Calibrating [271, 285]. calibration [57, 39, 89, 398]. Cancer [104, 168, 248, 87, 93, 250, 226, 213, 50, 247]. candidate [313]. candidate-set-free [313]. capture [317, 357]. capture-recapture [317]. carbon [55]. care [15, 47, 62]. Carlo [87, 114, 393, 39, 225, 90, 37, 308]. carotene [85]. carotid [388]. carrying [261]. Carter [120]. cartilage [151]. Case [109, 69, 157, 86, 267, 93, 286, 168, 44, 353, 11, 176, 259, 230]. Case-control [69, 93, 176]. Case-Study [109, 86, 267, 168, 353, 259, 230]. casting [119]. catastrophes [140]. catch [95, 149]. categorical [13, 283, 300]. cause [12, 117, 187]. cause-of-death [12, 187]. cell [354]. cement [333]. censored [394, 280, 159, 61, 365]. censoring [34, 310, 346]. census [349]. central [288]. centrelines [388]. cerebral [122]. chain [87, 114, 393, 39, 225, 90, 37, 308]. change [240, 199, 288, 73]. changed [58]. changepoint [58, 367, 38]. changepoints [159]. channel [241]. characteristic [332, 344, 65]. charges [62]. charts [177, 61]. Chebyshev [155]. checking [189]. checks [16]. chemicals [257]. Childhood [109, 14]. chlorophyll [285, 271]. chlorophyll-a [285, 271]. Choice [122, 318, 80]. chronologies [240, 355]. chronology [30]. City [164]. class [141, 142, 177, 27, 399, 83]. classification [253, 291, 99, 221, 307, 301]. classified [385]. climate [175, 398, 409]. clinical [233, 71, 182]. close [270]. cluster [298]. clustered [46, 364, 287, 365]. clustering [406]. coarsened [167]. code [154]. coefficients [139]. cognitive [73]. Cohort [107, 235, 63]. coin [190]. colour [314]. combinations [383]. Combining [169, 138, 338, 250, 226]. combustion [36]. common [311, 33]. comparing [232, 196]. comparison [46, 1, 59, 98, 139, 259]. comparisons [202]. Compartmental [132]. competing [117, 26, 61]. completely [173]. complex [302, 90, 336, 69]. compliance [306, 93, 79, 85]. complications [341]. component [183, 160]. components [395, 31]. Composite [131, 288, 6]. compositional [101, 362, 379]. Compound [105]. comprising [68]. computer [154]. concentrated [353]. concentrations [55]. concurrency [294]. Conditional [157, 13, 264, 33]. conditions [277]. confocal [151]. consequences [205]. consistency [217]. constraints [77, 268]. contact [270]. Contents [252, 331, 370, 414]. contests [265]. context [242]. contingency [326, 44]. continuous [204, 242, 333, 410, 402, 283, 300, 409, 259, 284]. control [177, 366, 93, 11, 176, 69, 61]. controls [45]. Conway [201]. Cooperative [86]. copula [383]. Cornwall [229]. correct [198, 259]. Correcting [93]. correction [175]. correlated [65, 375]. Correlating [242, 311]. Correlation [116, 233]. Correlation-Adjusted [116]. Correspondence [70, 20, 110]. Corrigendum [191, 403, 249, 379, 285, 59, 29, 250, 346, 300, 221]. cortex [277]. cosmic [319]. cost [393]. count [126, 400, 302, 363]. counterfactual [189, 267]. counting [234]. counts [255]. course [372]. covariance [224]. covariate [168, 317, 410, 84, 167, 209]. covariates [1, 59, 75, 11, 97]. Cox [75, 19]. crack [333]. credit [244]. Crime [8]. criterion [119]. crop [114, 80]. cross [203]. cross-design [203]. crossing [275]. crossover [94]. cumulative [263]. cure [86, 384]. current [335]. curtailment [121]. curve [399]. curves [316, 327, 65, 340, 388]. cut [178]. cycle [395]. cyclic [127]. cycling [184].

daily [100]. damage [365, 219]. Data [137, 138, 134, 201, 99, 101, 156, 233, 235, 318, 118, 381, 404, 126, 4, 212, 57, 362, 379, 157, 128, 10, 285, 271, 253, 195, 256, 325, 51, 279, 250, 226, 359, 287, 143, 60, 286, 278, 169, 299, 16, 202, 165, 397, 276, 400, 344, 168, 149, 79, 317, 67, 159, 302, 363, 304, 245, 123, 222, 84, 28, 41, 75, 274, 385, 19, 34, 8, 406, 335, 161, 37, 310, 346, 11, 282, 9, 83, 320, 100, 129, 311, 283, 61, 162, 365, 127, 259, 33, 68, 270, 221, 191, 403, 300]. data-driven [299]. database [244]. dated [355]. dating [77]. death [12, 187, 375]. deaths [213]. Debiased [377]. decline [181]. declines [368]. defect [199]. deficiency [128, 79, 375]. degree [180]. delays [358]. delineation [276]. demographic [118]. Denoising [102]. density [292]. dependence [58, 23]. dependences [126, 350]. dependent [115, 139, 210, 97]. depends [317]. deposition [343]. depression [212, 171]. depth [355]. Derbyshire [135]. derivative [236]. derivatives [388]. derived [53]. Design [130, 364, 220, 270, 119, 190, 288, 8, 203, 82]. designed [46]. Designing [170, 391, 6, 249]. Designs [314, 92, 185, 397, 158, 313, 304, 241]. desirability [21]. destination [72]. detectable [209]. Detecting [17, 266, 269]. Detection [175, 242, 319]. determination [9]. Development [103]. developmental [115, 173]. diagnoses [128]. diagnosing [368]. Diagnostic [16, 116, 290]. Diagnostics [348]. difference [153]. differences [259]. Different [138, 277, 47]. differential [195, 228, 269]. Differentiating [236]. diffusion [335]. dilution [359, 254]. dimensional [151, 327, 278, 302, 389, 388, 319]. dioxide [412]. diploid [408]. Direct [263]. directed [165]. disaggregation [146]. disclosure [326]. disconnected [158]. discoveries [246]. discovery [202]. Discrete [225, 201, 16, 202, 162]. discrimination [277]. discussion [262]. Disease [210, 215, 410, 123, 225, 37, 218, 347, 230, 270]. diseases [60, 14, 291]. dispersion [186]. display [302]. Distance [108]. distributed [372]. Distribution [108, 201, 234, 177, 323, 354, 38, 324, 296]. distribution-free [177]. Distributions [134, 394, 242, 52]. diversity [387, 193]. DNA [17, 389, 105]. dose [115, 281, 322, 208, 321, 80, 383, 411]. dose-response [281, 322, 321]. dosimeter [166]. Double [262, 28]. double-exponential [28]. downsizing [172]. draft [163]. driven [299]. drop [325, 79, 34, 218]. drop-out [325, 79]. drop-outs [34, 218]. dropout [363]. drug [43, 8, 307, 383]. drugs [232]. Dual [366, 241]. Duchenne [181]. due [24]. duration [339, 230]. durations [74]. during [277, 353]. dwell [406]. dwell-time-based [406]. Dynamic [106, 104, 268, 126, 47, 91, 182, 76, 49, 97, 179]. dynamical [14]. dynamics [22]. dystrophy [181].

E1690 [86]. Early [121]. earthquakes [147]. Eastern [86]. ecological [256]. Editors [371]. Effect [85, 103, 58, 1, 59, 143, 172, 64, 291]. Effectiveness [189]. effects [185, 267, 47, 287, 27, 194, 410, 64, 353, 360, 257, 181]. efficacy [306, 47, 411]. Efficient [272, 388, 118]. elderly [83]. electricity [211]. electroencephalographic [49]. emesis [233]. emotional [126]. emotions [238]. empirical [274, 309]. encephalopathy [63]. end [71, 350]. endogenous [342]. endoscopy [335]. engine [220]. England [120]. Environmental [106, 138, 240]. epidemic [215]. epidemiological [203]. Epidemiology [106]. epilepsy [122]. equation [84, 115, 282, 283, 300]. equations [2, 161]. equivalence [79]. error [152, 194, 84, 198, 209]. errors [119, 40, 239, 180, 264]. estimates [250, 226, 400, 54, 298]. Estimating [306, 342, 214, 267, 287, 149, 396, 43, 245, 408, 154, 254, 25, 235, 2, 84, 115, 337, 335, 161, 282, 372, 9, 171, 125, 409]. Estimation [60, 108, 116, 72, 180, 283, 12, 236, 144, 234, 392, 315, 292, 358, 199, 229, 143, 91, 187, 377, 96, 317, 67, 188, 89, 213, 336, 382, 208, 308, 388, 216, 230, 300]. Estimator [133]. estimators [339]. Europe [412, 98]. European [246]. evaluate [93, 353]. Evaluating [410]. Evaluation [156, 132, 351, 47, 393, 270, 191]. Event [136, 212, 255, 411]. events [18, 294, 375, 329, 68]. evidence [191, 156, 351, 203, 217]. Evolution [136]. exact [400]. example [64]. exchange [266]. exogenous [264]. expectancy [396]. expensive [407]. experiences [126]. experiment [46, 376, 94]. Experimental [304, 119, 277, 158, 222]. experimentation [220]. Experiments [130, 53, 314, 170, 249, 397, 215, 376, 302, 382, 309, 82]. exponential [21, 28]. exponentially [61]. exposure [353]. exposures [209]. Expression [137, 195, 228, 269]. Extended [359, 28]. extremal [18]. Extreme [260, 135, 140, 289, 405, 5, 35]. extremes [315, 206, 373].

facial [258, 238]. factor [250, 226, 268]. factorial [170, 249, 389]. Factors [32, 170, 15, 288, 53, 249]. failure [117, 71]. failures [74]. false [152, 202, 188]. familial [243]. families [10, 28]. family [173, 23, 243]. feature [238]. fertilizer [80]. fetal [402]. few [170, 249]. Fibre [134]. fidelity [220]. field [146, 166, 56, 276, 125]. fields [312, 6]. filaments [319]. finding [384, 383, 411]. finite [73, 298]. Finnish [33]. firing [277]. firm [345]. first [237, 335, 210]. first-order [237]. first-passage [210]. Fish [107]. fisheries [22]. fishery [95]. Fisz [299]. Fitting [201, 282]. flame [36]. Flexible [277, 54]. flow [279, 382, 48]. flows [405, 349]. fluid [48]. foetal [115]. Follicles [103]. follow [143]. follow-up [143]. food [253]. foraging [257]. force [287]. forces [145]. Forecasting [131, 120, 207]. forest [180]. forests [295]. form [191, 156]. formats [84]. fraction [188]. fractional [170, 249, 287]. Frailty [136]. free [318, 177, 313, 388]. free-choice [318]. free-knot [388]. frequencies [326]. frequency [266]. function [47, 263, 213, 181]. Functional [76, 280, 387, 153, 269, 341, 205]. Functions [108, 279, 21, 173]. future [246].

Gambia [109]. Garment [134]. gas [378]. gasifier [366]. Gaussian [379, 394, 146, 272, 4, 312, 362, 292, 52, 33, 6]. gear [53]. Geissler’s [28]. gene [195, 228, 269]. generalization [323, 324, 296]. Generalized [107, 206, 106, 167, 223, 341, 318, 56, 2, 84, 115, 65, 262, 336, 161, 282, 360, 120, 264]. generating [313]. generation [354]. Genetic [137, 248, 338, 284]. genetically [210]. genetics [169, 202]. genotype [169]. Geoadditive [112]. geochemistry [350]. geographical [198]. geological [88]. Geostatistics [109]. Gibbs [22]. GOES [299]. GOES-8 [299]. graphs [165]. greater [305]. group [377, 86, 168]. groups [169, 270]. growth [151, 399, 181]. gusts [35].

Haar [299]. hand [327]. handwriting [351]. haplotype [408]. haplotype-trait [408]. hard [366, 288]. hard-to-change [288]. hard-to-control [366]. harms [248]. harvesters [180]. Hastings [22]. hazard [406, 213]. hazards [26]. health [404, 396, 245, 198, 353, 360, 209, 62]. ‘Healthy [297]. heat [220, 360]. Heifers [103]. herd [287]. herd-specific [287]. heterogeneity [357, 216]. heterogeneous [147]. heteroscedastic [236]. heteroscedasticity [264, 62]. hidden [394, 17, 276, 396, 389, 33]. Hierarchical [138, 374, 219, 166, 391, 289, 193, 149, 79, 361, 123, 262, 372, 205]. High [278, 266, 302]. hip [219]. Hiroshima [213]. historical [45]. history [157]. HIV [128]. homogeneous [17, 276, 50, 246]. Hong [128]. Hormonal [103]. horticultural [282]. Horwitz [130]. hospital [15]. hospitalization [372]. household [37]. households [272]. human [235, 316, 189, 312, 128, 169, 202, 367, 43, 84, 28, 34, 218, 7, 97]. hurricane [18]. hurricanes [5]. hydraulic [53]. hydrocephalus [74]. hypothesis [239].

Identification [399]. identify [38]. Identifying [368]. ignorable [168, 96, 376, 363, 85]. ignorably [67]. ill [396]. image [151, 91, 400]. image-based [400]. images [114, 160, 179]. imaging [153, 76]. immune [128, 79, 84, 375]. immunity [230]. immunodeficiency [235, 189, 128, 169, 202, 367, 43, 84, 34, 218, 97]. immunology [46]. impact [256]. implants [115]. Improving [256]. imputation [297]. imputations [348]. incidence [235, 87, 128, 51, 187, 263, 123, 225, 54, 196, 217, 270]. income [144]. incompatible [84]. Incomplete [75, 96, 274, 19, 337, 406, 44]. incompletely [385]. Incorporating [269, 350]. Incorporation [45]. independence [239]. Index [42, 78, 111, 148, 192, 251, 293, 330, 369, 413, 131]. Indicators [131, 96]. individual [334, 157, 245, 200]. Industrial [131, 397]. inequality [155]. infant [267]. infarction [13]. infection [287, 410, 145, 25]. infectious [60, 37, 254, 230, 270]. Inference [276, 215, 56, 263, 40, 380, 227, 90, 320, 407]. inferences [208]. infinite [292]. Influence [107, 133, 290]. influencing [32]. influenza [230]. Information [131, 12, 187, 338]. informative [34, 310, 346, 218]. initial [306]. initiation [333]. injecting [43]. inner [388]. inspection [152]. integrating [6]. integration [118]. intensities [358, 400]. intensity [199]. interaction [194]. interactions [395]. interevent [38]. intermediately [54]. international [291, 168]. interobserver [198]. interval [365]. interval-censored [365]. interventions [270]. intestinal [261]. intraday [266]. investigate [7]. Investigation [135, 98]. irregularly [224]. issues [8]. item [290]. Iterative [102].

join [247]. Joint [255, 375, 68, 212, 365]. jointly [411]. Josephson [304]. jump [114, 393, 39, 308]. junctions [304].

Kalman [207]. karst [231]. kinematics [7]. knot [404, 388]. Kong [128]. kriged [207]. Kurtotic [134].

laboratory [197]. labour [125]. lag

[372, 49]. landmarks [160]. lapwings [368]. large [10, 244]. Latent [83, 146, 275, 362, 379, 26, 305, 333, 399, 343, 99, 221]. later [306]. lead [49]. learning [113]. leave [172]. Lee [120]. left [280, 34]. left-censored [280]. Leicester [72]. lethal [54]. lethality [12, 187]. level [373, 193]. levels [231, 164, 207]. life [396, 129, 297]. Likelihood [392, 234, 351, 96, 28, 274, 407]. limit [393]. Limits [197, 242]. line [294, 340, 259]. Linear [106, 185, 52, 194, 345, 262, 167, 22, 336, 268, 120, 181]. lines [48]. litter [173]. litter-based [173]. load [378]. Local [107, 358, 290, 125, 145]. local-influence-based [290]. location [223]. logistic [96, 64, 337, 198]. long [328, 54]. long-term [54]. longitudinal [233, 392, 212, 325, 79, 159, 363, 182, 19, 227, 90, 161, 310, 346, 218, 83, 259, 68, 301]. loss [143]. losses [295]. lottery [163]. Low [208, 144, 342, 180]. lower [335]. lung [181].

M. [135]. machine [334]. magnetic [166, 153, 76, 179]. major [353]. Malaria [109]. malignant [86]. management [113]. many [329, 94]. mapping [312]. maps [33]. marginal [52, 81]. marginalized [335]. marine [246]. mark [386]. mark-recapture-recovery [386]. marker [75]. markers [84, 218, 301]. market [149]. markets [125]. Markov [394, 146, 17, 87, 312, 280, 114, 56, 393, 39, 374, 396, 389, 225, 90, 37, 50, 308, 125]. masked [117]. Matched [11, 385]. maternal [63]. matrices [20]. Matrix [133, 391]. matter [372]. maximizing [21]. Maximum [91, 96, 234, 224]. Maxwell [201]. mean [288, 298]. measure [200]. measured [153]. measurement [258, 194, 84, 198, 180, 209]. measurements [81, 304, 23, 218]. measures [178, 88, 375, 68]. Measuring [345, 166, 291]. mechanical [21]. mechanisms [342]. Mechanistic [130]. melanoma [86]. member [350]. meningococcal [123]. mercury [343]. Meta [27, 119, 390, 24, 216, 338]. meta-analyses [338]. Meta-analysis [27, 390, 24, 216]. meta-model [119]. Metabolism [132]. method [214, 245, 274, 213, 282, 171]. methodology [157]. methods [87, 114, 393, 98, 39, 225, 353, 37, 203, 308, 99, 221]. metro [124]. Metropolis [22]. Mexico [164]. mice [261]. microarray [314, 155, 309, 241]. micronutrient [267]. microscopy [151]. minimizing [119]. Minimum [108]. misclassification [9]. Missing [97, 392, 168, 67, 376, 11, 83, 283, 300]. missingness [96]. mixed [185, 56, 278, 45, 340, 341, 181, 283, 300]. Mixture [195, 103, 292, 242, 178, 228, 297, 302, 384, 357]. mixtures [52, 406, 73, 25]. mobile [305]. Model [109, 104, 103, 132, 105, 298, 394, 146, 119, 295, 189, 404, 392, 4, 212, 378, 334, 362, 379, 280, 391, 292, 178, 195, 228, 297, 47, 289, 352, 290, 333, 149, 237, 396, 367, 164, 361, 363, 123, 288, 389, 75, 64, 124, 213, 335, 357, 227, 198, 90, 402, 350, 360, 372, 356, 211, 207, 181, 386, 283, 300, 247, 409, 55, 205, 35, 209, 217, 284, 174]. Model-Based [109, 298]. modelled [210]. Modelling [233, 1, 373, 279, 327, 193, 344, 138, 294, 147, 354, 343, 340, 257, 224, 181, 100, 145, 129, 347, 73, 5, 94, 394, 189, 258, 166, 277, 206, 140, 255, 359, 95, 14, 2, 305, 79, 84, 115, 28, 41, 8, 22, 225, 218, 120, 341, 219, 411, 59]. Models [130, 106, 136, 3, 159, 112, 185, 46, 126, 17, 141, 30, 86, 178, 256, 267, 287, 56, 278, 13, 26, 16, 27, 305, 215, 374, 96, 81, 194, 294, 302, 395, 122, 262, 399, 23, 182, 167, 384, 34, 239, 336, 180, 382, 45, 50, 49, 83, 340, 223, 341, 125, 264, 99, 127, 97, 29, 221]. modification [85]. modified [202]. Modifying [288]. molecule [222]. moment [214]. Monitoring [61, 292, 237, 6]. monotherapy [122]. monotone [287, 355, 173, 145]. monotonic [341]. monoxide [55]. Monte [87, 114, 393, 39, 225, 90, 37, 308]. mortality [87, 10, 267, 143, 120]. motion [327]. Motor [308]. mouse [29, 3]. movement [316]. moving [352, 61]. multi [349]. multi-resolution [349]. Multichannel [49]. multifaceted [403, 381]. multilevel [27, 174]. multimodal [279]. multinomial [178, 337]. multiple [71, 297, 202, 81, 188, 337, 161, 38, 100, 401]. multiple-changepoint [38]. multiple-comparisons [202]. multiple-screening [337]. multiple-source [161]. multiple-testing [401]. multiplicity [29, 3, 261]. multiranked [124]. multiscale [299]. multistate [235, 365]. multistratum [397]. Multivariate [232, 406, 348, 191, 156, 4, 10, 229, 278, 52, 276, 399, 100, 55, 284, 6]. muscular [181]. myocardial [13].

natural [378]. Near [241]. Near-optimal [241]. needle [295]. negative [392, 188]. negatives [188]. neglog [244]. networks [165, 6]. neural [277]. neuron [277]. no [152]. noise [236, 36]. noisy [236]. Non [231, 52, 168, 317, 22, 50, 306, 93, 373, 397, 344, 79, 96, 67, 376, 363, 85, 345, 41, 399, 408, 384, 268, 209, 339, 246, 298]. non-compliance [306, 93, 79, 85]. non-detectable [209]. non-diploid [408]. Non-homogeneous [50, 246]. Non-ignorable [168, 96, 376, 363, 85]. non-ignorably [67]. Non-linear [52, 22, 268]. non-mixture [384]. non-normal [41]. non-orthogonal [397]. Non-parametric [317, 344, 345, 339]. non-responders [399]. non-response [298]. Non-stationary [231, 373].

Nonparametric [234, 199, 229, 88, 89, 213, 155, 245, 65, 227, 38]. normal [41, 341]. north [95, 315]. north-west [95]. novel [142, 402]. nuclear [186]. number [60, 172, 115, 64, 308]. numbers [254].

object [319]. objectives [325]. observed [410, 224]. Obtaining [226, 250]. occult [12]. occupational [209]. Occurrences [105, 147]. oceanographic [4]. olfactometer [320]. on-line [294, 340]. Oncology [86, 383]. onset [12]. operating [332, 344, 65]. opinion [403, 381]. Optimal [176, 66, 113, 313, 82, 329, 241]. optimization [366, 21]. oral [198]. order [344, 237, 77, 281, 322, 321]. Ordered [108, 283, 300, 99, 221]. ordinal [81, 182, 198, 282, 94]. orientation [98, 7]. origin [72]. orthogonal [318, 397]. orthopaedic [219]. oscillation [214]. outbreaks [37]. outcome [24]. outcomes [265, 46, 392, 67, 376, 402, 356, 283, 300, 411]. outliers [297]. outputs [154]. outs [34, 218]. Overcoming [401]. overdispersion [41]. owing [61]. ozone [373, 164].

paediatric [79, 74]. paired [1, 59, 98]. pairs [385]. palsy [122]. panel [283, 300]. papilloma [29, 3]. parallel [120]. parameter [67]. parameters [214, 9, 407, 409]. Parametric [104, 384, 344, 317, 263, 122, 345, 291, 339, 303]. partial [98]. partially [117, 99, 221]. particle [382]. particles [305, 353]. particulate [372]. partitioning [169]. passage [210]. passenger [124]. paths [340]. pathway [184]. patient [143]. patients [74]. pattern [334, 172, 36]. patterns [90, 353]. pay [342]. peat [240]. penalized [285, 271]. percentile [267]. percentile-specific [267]. percentiles [154]. perception [238]. performance [366, 345, 337]. Period [150, 328, 94]. periodicities [266]. periods [328]. permanent [391]. phage [302]. pharmaceutical [152]. pharmacokinetic [66, 82]. phases [38]. photometric [361]. phylogenetic [389]. physiological [153]. pill [152]. pilot [158]. pine [295]. pine-trees [295]. pipe [48]. pipe-lines [48]. plant [347]. plants [25]. plebiscite [44]. plot [170, 249, 390, 313]. plots [391]. point [178, 229, 88, 311, 319, 247]. point-referenced [311]. points [71, 64]. Poisson [358, 359, 261, 105, 201]. policy [113]. pollutant [100]. pollution [92, 353, 372, 207, 6]. polymorphisms [205]. polynomial [358, 180]. polynomials [287, 145]. Population [393, 234, 391, 317, 368, 66, 298]. Population-based [393]. populations [40, 408]. positive [128, 152]. posterior [16]. posteriori [91]. potato [25]. potent [189]. power [173, 139]. precipitation [394]. Predicting [265, 246, 326, 374, 205]. prediction [119, 400, 180, 55]. predictive [16, 176]. preference [245]. prefrontal [277]. Prepubertal [103]. presence [204, 236, 280, 358, 332, 297, 375, 243, 218]. prevalence [250, 226, 187]. prevalent [235]. prevention [248]. price [340]. prices [211]. primary [15]. Principal [160, 31]. priors [56]. Probabilistic [351, 238]. probabilities [232, 342]. Probability [292, 43, 317, 33]. probit [124]. problem [38, 401]. procedure [142, 202, 309]. procedures [158, 259]. process [119, 280, 292, 256, 229, 359, 88, 355, 220, 288, 311, 319, 33, 246]. process-based [256]. processes [375, 343, 210, 224, 61, 365]. Procrustes [318]. Production [131, 152]. products [47]. profile [318]. profiles [212, 378, 387]. programmes [93]. programming [91]. progression [189, 218]. project [297]. projecting [87]. projections [120]. prone [5]. propagation [25]. properties [21, 398]. prophylactic [270]. proportion [144]. Proportional [26, 406]. proportions [377]. proposal [325]. proposed [248, 306]. psoriatic [365]. Psychotropic [307]. public [403, 381]. pulsation [328]. pumps [53].

quadrature [272, 64]. quantal [196]. quantile [345, 244]. quantitation [197]. quasi [28]. quasi-likelihood [28]. questionnaire [278, 8].

radar [57]. radial [390]. radiation [352]. Radiocarbon [240, 77, 355]. radiocarbon-dated [355]. Radiocommunications [102, 236]. radon [311]. rainfall [146, 57]. Random [108, 34, 146, 312, 47, 287, 56, 88, 64, 83, 125]. random-effects [47, 287]. random-field [146, 125]. Randomization [137, 303]. randomized [248, 306, 71, 376, 85]. randomness [401]. rankings [1, 59]. ranks [98]. rare [40, 239]. Rate [136, 86, 202, 25]. rates [266, 277, 15, 72, 196, 247]. ratio [351, 28]. rats [307]. ray [299]. Realtime [407]. recalled [171]. recapture [317, 386]. Receiver [332, 344, 65]. recognition [36]. Reconstructing [128]. recovery [47, 386]. Recurrent [136, 375]. recursive [169]. red [361]. referenced [311]. regions [276, 153, 5]. registry [143]. Regression [133, 2, 188, 127, 62, 295, 392, 285, 271, 29, 3, 279, 16, 305, 237, 337, 34, 198, 402, 49, 388, 247, 244, 383, 284]. Regression-based [188]. regressions [345, 25]. regressive [352]. related [214, 227, 338]. relationship [142, 28]. relative [30, 200]. reliability [166, 219]. remotely [285, 271]. removal [404]. renewal [246]. Repeated [152, 178, 74, 23, 375, 356, 68]. repeatedly [356]. repellent [257]. replacement [219]. replicates [189]. replication [309]. Report [371]. reporting [358]. reproduction [60, 380]. research [155, 282]. residual [147]. resistance [210]. resolution [349]. resonance [153, 76, 179]. resource [364]. resources [380]. respiratory [404]. responders [399]. Response [303, 366, 273, 13, 290, 115, 153, 182, 196, 281, 322, 324, 321, 94, 298]. Response-adaptive [303]. responses [96, 274, 19]. restriction [281, 322, 321]. restrictions [393, 406]. results [152]. reuse [82]. reversible [114, 393, 39, 308, 160]. revision [291]. revisited [63, 163]. Revival [201]. ribonucleic [34]. risk [250, 226, 326, 26, 405, 173, 372, 208]. risks [117, 212, 61]. river [405]. Robertsonian [261]. Robust [185, 237, 150]. Rule [130]. rules [253].

safety [232]. Sample [82, 275, 391, 206, 367, 65, 9, 176]. sampled [340]. Samples [137, 298]. sampling [92, 272, 149, 22, 336, 69, 66]. scale [223, 94]. Scales [138]. schedule [384]. score [282]. scoring [244]. screening [204, 93, 152, 337]. Sea [315]. Seasonality [395]. section [151]. sediment [350]. segment [58]. segmentation [114]. Segmenting [389]. segments [17]. segregation [229]. seismicity [38]. selection [342, 364, 393, 24, 213, 310, 346, 360, 386, 264, 247]. semi [280]. semi-Markov [280]. semicontinuous [233]. Semiparametric [178, 286, 305, 367, 230, 301, 294, 45, 356, 340, 174]. sensed [285, 271]. Sensitivity [116, 44, 403, 381]. sensor [299]. sequence [58, 389]. Sequences [105, 17]. sequential [190, 121, 407]. serial [233, 23, 254]. series [141, 175, 359, 286, 14, 52, 237, 225, 55, 230]. serological [60]. session [406]. set [88, 313]. sex [28]. sexual [380]. shape [260, 223]. sheep [32]. shifts [175, 361]. short [207]. short-term [207]. shovelnose [374]. Shrinkage [102]. shunt [74]. sick [172]. Sides [130]. signalling [184]. Signals [102, 236]. Simple [158, 183, 216, 31, 339, 355, 281, 322, 321, 409].

Simulation [18, 82]. simulations [16]. Simultaneous [21, 208, 329]. single [222]. single-molecule [222]. Sinking [135]. sites [100]. size [234, 142, 27, 317, 9, 176]. sizes [40, 243]. skin [47]. sleep [307]. sleep-wake [307]. Slovenian [403, 381, 44]. Small [15, 137, 155, 250, 226, 326]. smoothed [98]. smoothing [412, 150, 76, 340, 409]. Soay [32]. social [165]. solar [352]. Solving [113]. source [305, 161]. sources [353]. Space [394, 57, 123, 22, 147, 343, 347]. Space-time [394, 57, 123, 343]. spacing [237]. sparse [340]. Spatial [400, 405, 92, 295, 229, 88, 276, 333, 163, 350, 311]. spatially [391]. Spatiotemporal [412, 146, 231, 352, 305, 164, 294, 224]. spawning [374]. species [193, 246]. specific [1, 59, 267, 287]. Specificity [116]. Spectral [36]. speed [279]. speed-flow [279]. speeded [290]. speeds [289, 5]. spells [280]. spline [404]. splines [271, 388, 285]. split [170, 249, 313]. split-plot [170, 249, 313]. spongiform [63]. sporting [265]. spread [347]. square [20]. Stability [404]. stabilization [299]. stable [354]. staff [172]. stage [298]. standardized [378]. stars [328, 150]. state [126, 22]. static [407]. stationary [231, 373]. Statistical [151, 376, 353, 320, 398, 378, 334, 113, 292, 256, 350, 208]. Statistics [248, 133, 65, 7, 349]. status [280, 410, 335]. steel [21]. stellar [214]. stellar-oscillation-related [214]. Stimulation [103]. stochastic [344, 215, 121, 210, 409]. Stochastically [108]. Stock [107]. stopping [121]. stratified [60]. structure [49]. structured [295]. studies [232, 275, 364, 253, 1, 59, 93, 29, 3, 143, 367, 24, 115, 173, 23, 227, 402, 281, 322, 83, 176, 69, 321, 66, 338, 241]. Study [109, 235, 212, 86, 63, 267, 27, 168, 85, 375, 198, 353, 259, 209, 230]. sturgeon [374]. subject [242, 1, 59, 34, 310, 346, 9]. subject-specific [1, 59]. subjective [94]. subset [264]. substance [310, 346]. substudy [248]. success [374]. sulphur [412]. superconducting [304]. supplementation [267]. surface [366, 373]. surfaces [332]. surges [315]. surrogate [71]. surveillance [123]. survey [272, 403, 381, 60, 274, 336, 8]. surveys [226, 250]. Survival [104, 32, 255, 168, 122, 196, 50, 200, 174, 303]. survival-adjusted [196]. survivors [213]. sustainable [366]. switching [374, 304]. syndrome [128, 79, 375]. synthesis [256, 203, 217]. synthetic [315]. system [398]. systems [14].

table [385]. tables [326, 44]. tail [4]. task [277]. teller [334]. tempered [354]. Temporal [172, 334, 77, 268]. term [54, 207]. tessellation [276]. test [204, 275, 337, 196, 129]. testing [155, 377, 408, 401]. Tests [137, 116, 153, 184, 281, 322, 321, 232, 58, 128]. texture [114]. Thailand [43]. their [214, 388]. theory [290]. therapeutic [384]. therapy [189]. there [288]. three [327, 121, 388, 319]. three-arm [121]. three-dimensional [327, 388, 319]. threshold [275, 211]. thresholding [309]. Time [136, 14, 12, 394, 141, 57, 71, 52, 237, 123, 406, 225, 147, 372, 38, 49, 343, 347, 162, 409, 55, 97, 230, 411]. time-dependent [97]. time-to-event [411]. time-varying [49]. Times [104, 255, 354, 210, 66, 68]. timing [153]. tissue [341]. Tmax [82]. toad [33]. tocopherol [85]. toxicity [115, 173, 402, 411]. toxicological [203]. toxoplasmosis [217]. trace [191, 156]. Tracing [143]. traffic [141]. trait [408]. traits [284]. trajectories [327]. trans [389]. trans-dimensional [389]. transform [299]. transformation [19, 244]. Transforming [130]. transgenic [29, 3]. transition [342]. transitions [280]. translocation [261]. transmission [215, 43]. treatment [273, 267, 13, 220, 310, 346]. treatments [190]. tree [281, 322, 321]. trees [295]. Trend [315, 395]. trends [412]. trial [233, 248, 306, 86, 168, 79, 85, 121, 182]. trials [71, 303]. trip [72]. tuberculosis [229]. tumorigenicity [51]. Tumour [187, 12, 273, 261, 54, 196]. tumours [12, 54]. turbulent [349]. Two [116, 41, 151, 190, 314, 275, 166, 242, 250, 226, 367, 153, 65, 218, 224, 283, 300, 298]. two-colour [314]. two-dimensional [151]. two-equation [283, 300]. two-sample [275, 367, 65]. two-stage [298]. type [169, 218]. types [166]. typical [54].

UK [229, 368, 217]. uncertain [154]. uncertainties [214]. understand [380]. understanding [73]. unemployment [339]. unequal [237, 40]. unevenly [340]. unified [209]. uniform [23]. unit [40, 308, 298]. units [254]. university [1, 59]. unlabelled [253]. unmatched [364]. unobserved [395]. unverified [188]. update [253]. updating [186]. uranium [311]. urban [92]. use [335, 94]. Useful [201]. users [43]. Using [102, 253, 297, 67, 385, 19, 380, 7, 394, 235, 404, 17, 280, 391, 128, 285, 271, 256, 114, 287, 327, 98, 299, 16, 27, 2, 149, 215, 396, 410, 345, 153, 399, 121, 274, 182, 213, 22, 225, 161, 354, 45, 360, 76, 340, 308, 224, 145, 129, 61, 33, 246, 270]. utilities [245].

V [135]. Validation [71]. valuation [275]. Value [135, 140, 98]. values [385]. valve [220]. variable [393, 305, 328, 220, 24, 150, 310, 346]. Variables [108, 275, 242, 366, 264]. Variance [144, 336, 195, 299, 288, 216]. variation [15, 261]. variations [7]. variety [80]. varying [243, 49]. vector [76]. verification [204, 332, 337]. via [272, 189, 113, 292, 294, 173, 124, 208, 49, 25]. viral [389, 25]. virus [235, 189, 128, 169, 202, 367, 43, 84, 34, 97]. volume [233, 42, 78, 111, 148, 192, 252, 251, 293, 331, 330, 370, 369, 414, 413, 141]. volumetric [39]. vortex [349]. voting [376].

wake [307]. Wales [120]. warranty [162]. water [231, 237]. wave [360]. Wavelet [102, 236]. ways [41]. weed [114, 400]. Weibull [406]. weighing [248]. weight [267]. weighted [2, 61]. weights [76]. west [95]. while [209]. whole [170, 249]. whole-plot [170, 249]. wholesale [211]. Wilcoxon [65]. wind [289, 5, 35]. withdrawals [334]. within [24, 22]. within-Gibbs [22]. Word [105]. Württemberg [295].

X [299]. X-ray [299].

yield [113].

zeros [362, 379].

References

Dittrich:1998:MES

[1] R. Dittrich, R. Hatzinger, and W. Katzenbeisser. Modelling the effect of subject-specific covariates in paired comparison studies with an application to university rankings. Applied Statistics, 47(4):511–??, 1998. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

                                             http://www.blackwellpublishers.
                                             co.uk/asp/journal.asp?ref=0035-
                                             9254&src=ard&aid=125&iid=4&vid=

47. See corrigendum [59].

Gonin:2000:RMW

[2] R. Gonin, S. R. Lipsitz, G. M. Fitzmaurice, and G. Molenberghs. Regression modelling of weighted κ by using generalized estimating equations.

Applied Statistics, 49(1):1–18, 2000. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=175&iid=1&vid=
49.

Dunson:2000:MPM

[3] D. B. Dunson. Models for papilloma multiplicity and regression: applications to transgenic mouse studies. Applied Statistics, 49(1):19–30, 2000. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=176&iid=1&vid=

49. See corrigendum [29].

Bortot:2000:MGT

[4] P. Bortot, S. Coles, and J. Tawn. The multivariate Gaussian tail model: an application to oceanographic data. Applied Statistics, 49(1):31–49, 2000. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=177&iid=1&vid=
49.

Walshaw:2000:MEWa

[5] D. Walshaw. Modelling extreme wind speeds in regions prone to hurricanes. Applied Statistics, 49(1):51–62, 2000. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=178&iid=1&vid=

Zidek:2000:DIC

[6] J. V. Zidek, W. Sun, and N. D. Le. Designing and integrating composite networks for monitoring multivariate Gaussian pollution fields. Applied Statistics, 49(1):63–79, 2000. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

                                        http://www.blackwellpublishers.
                                        co.uk/asp/journal.asp?ref=0035-
                                        9254&src=ard&aid=179&iid=1&vid=
                                        49.

Rancourt:2000:UOS

[7] D. Rancourt, L.-P. Rivest, and J. Asselin. Using orientation statistics to investigate variations in human kinematics. Applied Statistics, 49(1):81–94, 2000. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http:

                                        //www.blackwellpublishers.co.uk/
                                        asp/journal.asp?ref=0035-9254&src=
                                        ard&aid=180&iid=1&vid=49.

MacDonald:2000:ADA

[8] Z. MacDonald and S. Pudney. Analysing drug abuse with British Crime Survey data: modelling and questionnaire design issues. Applied Statistics, 49(1):95–117, 2000. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

                                        http://www.blackwellpublishers.
                                        co.uk/asp/journal.asp?ref=0035-
                                        9254&src=ard&aid=181&iid=1&vid=
                                        49.

Rahme:2000:BSS

[9] E. Rahme, L. Joseph, and T. W. Gyorkos. Bayesian sample size determination for estimating binomial parameters from data subject to misclassification.

Applied Statistics, 49(1):119–128, 2000. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=182&iid=1&vid=
49.

Chen:2000:BAM

[10] M.-H. Chen, D. K. Dey, and D. Sinha. Bayesian analysis of multivariate mortality data with large families. Applied Statistics, 49(1):129–144, 2000. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=183&iid=1&vid=
49.

Paik:2000:MCC

[11] M. C. Paik and R. L. Sacco. Matched case–control data analyses with missing covariates. Applied Statistics, 49(1):145–156, 2000. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=184&iid=1&vid=
49.

Ahn:2000:ATL

[12] H. Ahn, H. Moon, and R. L. Kodell. Attribution of tumour lethality and estimation of the time to onset of occult tumours in the absence of cause-of-death information. Applied Statistics, 49(2):157–169, 2000. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-

                                        9254&src=ard&aid=185&iid=2&vid=
                                        49.

Gelfand:2000:CCR

[13] A. E. Gelfand, M. D. Ecker, C. Christiansen, T. J. McLaughlin, and S. B. Soumerai. Conditional categorical response models with application to treatment of acute myocardial infarction. Applied Statistics, 49(2):171–186, 2000. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

                                        http://www.blackwellpublishers.
                                        co.uk/asp/journal.asp?ref=0035-
                                        9254&src=ard&aid=186&iid=2&vid=
                                        49.

Finkenstadt:2000:TSM

[14] B. F. Finkenstädt and B. T. Grenfell. Time series modelling of childhood diseases: a dynamical systems approach. Applied Statistics, 49(2):187–205, 2000. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

                                        http://www.blackwellpublishers.
                                        co.uk/asp/journal.asp?ref=0035-
                                        9254&src=ard&aid=187&iid=2&vid=
                                        49.

Congdon:2000:SAV

[15] P. Congdon and N. Best. Small area variation in hospital admission rates: Bayesian adjustment for primary care and hospital factors. Applied Statistics, 49(2):207–226, 2000. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

                                        http://www.blackwellpublishers.
                                        co.uk/asp/journal.asp?ref=0035-
                                        9254&src=ard&aid=188&iid=2&vid=
                                        49.

Gelman:2000:DCD

[16] A. Gelman, Y. Goegebeur, F. Tuerlinckx, and I. Van Mechelen. Diagnostic checks for discrete data regression models using posterior predictive simulations. Applied Statistics, 49(2):247–268, 2000. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=190&iid=2&vid=
49.

Boys:2000:DHS

[17] R. J. Boys, D. A. Henderson, and D. J. Wilkinson. Detecting homogeneous segments in DNA sequences by using hidden Markov models. Applied Statistics, 49(2):269–285, 2000. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=191&iid=2&vid=
49.

Casson:2000:SEA

[18] E. Casson and S. Coles. Simulation and extremal analysis of hurricane events. Applied Statistics, 49(3):227–245, 2000. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=189&iid=3&vid=
49.

Lipsitz:2000:UBC

[19] S. R. Lipsitz, J. Ibrahim, and G. Molenberghs. Using a Box–Cox transformation in the analysis of longitudinal data with incomplete responses. Applied Statistics, 49(3):287–296, 2000. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=192&iid=3&vid=
49.

Greenacre:2000:CAS

[20] Michael Greenacre. Correspondence analysis of square asymmetric matrices. Applied Statistics, 49(3):297–310, 2000. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=193&iid=3&vid=
49.

Kim:2000:SOM

[21] Kwang-Jae Kim and Dennis K. J. Lin. Simultaneous optimization of mechanical properties of steel by maximizing exponential desirability functions. Applied Statistics, 49(3):311–325, 2000. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=194&iid=3&vid=
49.

Millar:2000:NLS

[22] Russell B. Millar and Renate Meyer. Non-linear state space modelling of fisheries biomass dynamics by using Metropolis–Hastings within-Gibbs sampling. Applied Statistics, 49 (3):327–342, 2000. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http:

//www.blackwellpublishers.co.uk/
asp/journal.asp?ref=0035-9254&src=
ard&aid=195&iid=3&vid=49.

Lindsey:2000:FMU

[23] J. K. Lindsey. A family of models for uniform and serial dependence in repeated measurements studies. Applied Statistics, 49(3):343–357, 2000. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=196&iid=3&vid=
49.

Hutton:2000:BMA

[24] J. L. Hutton and Paula R. Williamson. Bias in meta-analysis due to outcome variable selection within studies. Applied Statistics, 49(3):359–370, 2000. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=197&iid=3&vid=
49.

Turner:2000:EPR

[25] T. Rolf Turner. Estimating the propagation rate of a viral infection of potato plants via mixtures of regressions. Applied Statistics, 49(3):371–384, 2000. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=198&iid=3&vid=
49.

Gelfand:2000:PHM

[26] Alan E. Gelfand, Sujit K. Ghosh, Cindy Christiansen, Stephen B. Soumerai, and Thomas J. McLaughlin. Proportional hazards models: a latent competing risk approach. Applied Statistics, 49(3):385–397, 2000. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=199&iid=3&vid=
49.

Goldstein:2000:MAU

[27] Harvey Goldstein, Min Yang, Rumana Omar, Rebecca Turner, and Simon Thompson. Meta-analysis using multilevel models with an application to the study of class size effects. Applied Statistics, 49(3):399–412, 2000. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=200&iid=3&vid=
49.

Lee:2000:RBD

[28] Y. Lee and J. A. Nelder. The relationship between double-exponential families and extended quasi-likelihood families, with application to modelling Geissler’s human sex ratio data. Applied Statistics, 49(3):413–419, 2000. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=201&iid=3&vid=
49.

Dunson:2000:CMP

[29] D. B. Dunson. Corrigendum: Models for papilloma multiplicity and regression: applications to transgenic mouse studies. Applied Statistics, 49(3):421–422, 2000. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.

co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=202&iid=3&vid=

49. See [3].

Buck:2000:BMR

[30] Caitlin E. Buck and Sujit K. Sahu. Bayesian models for relative archaeological chronology building. Applied Statistics, 49(4):423–440, 2000. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=203&iid=4&vid=
49.

Vines:2000:SPC

[31] S. K. Vines. Simple principal components. Applied Statistics, 49 (4):441–451, 2000. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http:

//www.blackwellpublishers.co.uk/
asp/journal.asp?ref=0035-9254&src=
ard&aid=204&iid=4&vid=49.

Catchpole:2000:FIS

[32] E. A. Catchpole, B. J. T. Morgan, T. N. Coulson, S. N. Freeman, and S. D. Albon. Factors influencing Soay sheep survival. Applied Statistics, 49(4):453–472, 2000. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=205&iid=4&vid=
49.

Weir:2000:BPM

[33] I. S. Weir and A. N. Pettitt. Binary probability maps using a hidden conditional autoregressive Gaussian process with an application to Finnish common toad data. Applied Statistics, 49(4):473–484, 2000. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=206&iid=4&vid=
49.

Lyles:2000:RRM

[34] Robert H. Lyles, Cynthia M. Lyles, and Douglas J. Taylor. Random regression models for human immunodeficiency virus ribonucleic acid data subject to left censoring and informative drop-outs. Applied Statistics, 49(4):485–497, 2000. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=207&iid=4&vid=
49.

Walshaw:2000:MEWb

[35] David Walshaw and Clive W. Anderson. A model for extreme wind gusts. Applied Statistics, 49(4):499–508, 2000. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=208&iid=4&vid=
49.

Yuan:2000:SAC

[36] J. Yuan and Y. Zhang. Spectral analysis of combustion noise and flame pattern recognition. Applied Statistics, 49(4):509–515, 2000. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-

9254&src=ard&aid=209&iid=4&vid=
49.

ONeill:2000:AID

[37] Philip D. O’Neill, David J. Balding, Niels G. Becker, Mervi Eerola, and Denis Mollison. Analyses of infectious disease data from household outbreaks by Markov chain Monte Carlo methods. Applied Statistics, 49(4):517–542, 2000. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=210&iid=4&vid=
49.

Pievatolo:2000:AIT

[38] Antonio Pievatolo and Renata Rotondi. Analysing the interevent time distribution to identify seismicity phases: a Bayesian nonparametric approach to the multiple-changepoint problem. Applied Statistics, 49(4):543–562, 2000. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=211&iid=4&vid=
49.

Henderson:2000:BVC

[39] Robin Henderson, Tony Morton-Jones, and Paul McKnespiey. Bayesian volumetric calibration: an application of reversible jump Markov chain Monte Carlo methods. Applied Statistics, 49(4):563–575, 2000. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.
co.uk/asp/journal.asp?ref=0035-
9254&src=ard&aid=212&iid=4&vid=

Laws:2000:BIR

[40] David J. Laws and Anthony O’Hagan. Bayesian inference for rare errors in populations with unequal unit sizes. Applied Statistics, 49(4):577–590, 2000. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

                                        http://www.blackwellpublishers.
                                        co.uk/asp/journal.asp?ref=0035-
                                        9254&src=ard&aid=213&iid=4&vid=
                                        49.

Lee:2000:TWM

[41] Y. Lee and J. A. Nelder. Two ways of modelling overdispersion in non-normal data. Applied Statistics, 49(4):591–598, 2000. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

                                        http://www.blackwellpublishers.
                                        co.uk/asp/journal.asp?ref=0035-
                                        9254&src=ard&aid=214&iid=4&vid=
                                        49.

Anonymous:2000:IAV

[42] Anonymous. Index of authors, volume 49, 2000. Applied Statistics, 49(4):??, 2000. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http:

                                        //www.blackwellpublishers.co.uk/
                                        asp/journal.asp?ref=0035-9254&src=
                                        ard&aid=215&iid=4&vid=49.

Hudgens:2001:ETP

[43] Michael G. Hudgens, Ira. M. Longini, Jr., M. Elizabeth Halloran, Kachit Choopanya, Suphak Vanichseni, Dwip Kitayaporn, Timothy D. Mastro, and Philip A. Mock. Estimating the transmission probability of human immunodeficiency virus in injecting drug users in Thailand. Applied Statistics, 50(1):1–??, 2001. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Molenberghs:2001:SAI

[44] Geert Molenberghs, Michael G. Kenward, and Els Goetghebeur. Sensitivity analysis for incomplete contingency tables: the Slovenian plebiscite case. Applied Statistics, 50(1):15–??, 2001. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Parise:2001:IHC

[45] Helen Parise, M. P. Wand, David Ruppert, and Louise Ryan. Incorporation of historical controls using semiparametric mixed models. Applied Statistics, 50(1):31–??, 2001. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Betensky:2001:CMC

[46] Rebecca A. Betensky, Paige L. Williams, and Howard M. Lederman. A comparison of models for clustered binary outcomes: analysis of a designed immunology experiment. Applied Statistics, 50(1):43–??, 2001. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Erlandsen:2001:DRE

[47] Mogens Erlandsen and Lars Halkier-Sørensen. A dynamic random-effects model for recovery of skin barrier function: evaluation of the efficacy of different skin care products. Applied Statistics, 50(1):63–??, 2001. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Rougier:2001:BAF

[48] Jonathan Rougier and Michael Goldstein. A Bayesian analysis of fluid flow in pipe-lines. Applied Statistics, 50(1):77– ??, 2001. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Prado:2001:MEA

[49] Raquel Prado, Mike West, and Andrew D. Krystal. Multichannel electroencephalographic analyses via dynamic regression models with timevarying lag–lead structure. Applied Statistics, 50(1):95–??, 2001. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Perez-Ocon:2001:NHM

[50] Rafael Pérez-Ocón, Juan Eloy Ruiz-Castro, and M. Luz Gámiz-Pérez. Nonhomogeneous Markov models in the analysis of survival after breast cancer. Applied Statistics, 50(1):111–??, 2001. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Dunson:2001:BIA

[51] D. B. Dunson and G. E. Dinse. Bayesian incidence analysis of animal tumorigenicity data. Applied Statistics, 50(2): ??, 2001. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Glasbey:2001:NLA

[52] C. A. Glasbey. Non-linear autoregressive time series with multivariate Gaussian mixtures as marginal distributions. Applied Statistics, 50(2):??, 2001. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Sexton:2001:EDF

[53] C. J. Sexton, S. M. Lewis, and C. P. Please. Experiments for derived factors with application to hydraulic gear pumps. Applied Statistics, 50(2):??, 2001. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Parise:2001:FET

[54] Helen Parise, Gregg E. Dinse, and Louise M. Ryan. Flexible estimates of tumour incidence for intermediately lethal tumours in a typical long-term animal bioassay. Applied Statistics, 50(2): ??, 2001. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Tonellato:2001:MTS

[55] Stefano F. Tonellato. A multivariate time series model for the analysis and prediction of carbon monoxide atmospheric concentrations. Applied Statistics, 50(2):??, 2001. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Fahrmeir:2001:BIG

[56] Ludwig Fahrmeir and Stefan Lang. Bayesian inference for generalized additive mixed models based on Markov random field priors. Applied Statistics, 50(2):??, 2001. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Brown:2001:STC

[57] Patrick E. Brown, Peter J. Diggle, Martin E. Lord, and Peter C. Young. Spacetime calibration of radar rainfall data. Applied Statistics, 50(2):??, 2001. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Avery:2001:EDB

[58] Peter J. Avery. The effect of dependence in a binary sequence on tests for a changepoint or a changed segment. Applied Statistics, 50(2):??, 2001. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Dittrich:2001:CME

[59] R. Dittrich, R. Hatzinger, and W. Katzenbeisser. Corrigendum: Modelling the effect of subject-specific covariates in paired comparison studies with an application to university rankings. Applied Statistics, 50(2):??, 2001. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See [1].

Farrington:2001:EBR

[60] C. P. Farrington, M. N. Kanaan, and N. J. Gay. Estimation of the basic reproduction number for infectious diseases from age-stratified serological survey data. Applied Statistics, 50(3): ??, ???? 2001. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Steiner:2001:MPD

[61] Stefan H. Steiner and R. Jock MacKay. Monitoring processes with data censored owing to competing risks by using exponentially weighted moving average control charts. Applied Statistics, 50(3): ??, ???? 2001. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Zhou:2001:RAH

[62] Xiao-Hua Zhou, Kevin T. Stroupe, and William M. Tierney. Regression analysis of health care charges with heteroscedasticity. Applied Statistics, 50(3):??, ???? 2001. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Cooper:2001:ABS

[63] J. D. Cooper and S. M. Bird. Analysis of the bovine spongiform encephalopathy maternal cohort study, revisited. Applied Statistics, 50(3):??, ???? 2001. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lesaffre:2001:ENQ

[64] Emmanuel Lesaffre and Bart Spiessens. On the effect of the number of quadrature points in a logistic random effects model: an example. Applied Statistics, 50(3):??, ???? 2001. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lee:2001:AAU

[65] Mei-Ling Ting Lee and Bernard A. Rosner. The average area under correlated receiver operating characteristic curves: a nonparametric approach based on generalized two-sample Wilcoxon statistics. Applied Statistics, 50(3):??, ???? 2001. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Stroud:2001:OST

[66] Jonathan R. Stroud, Peter Müller, and Gary L. Rosner. Optimal sampling times in population pharmacokinetic studies. Applied Statistics, 50(3):??, ???? 2001. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ibrahim:2001:UAD

[67] Joseph G. Ibrahim, Stuart R. Lipsitz, and Nick Horton. Using auxiliary data for parameter estimation with nonignorably missing outcomes. Applied Statistics, 50(3):??, ???? 2001. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Xu:2001:JAL

[68] Jane Xu and Scott L. Zeger. Joint analysis of longitudinal data comprising repeated measures and times to events. Applied Statistics, 50(3):??, ???? 2001. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Scott:2001:CCS

[69] Alastair Scott and Chris Wild. Casecontrol studies with complex sampling. Applied Statistics, 50(3):??, ???? 2001. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:2001:C

[70] Anonymous. Correspondence. Applied Statistics, 50(3):??, ???? 2001. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Burzykowski:2001:VSE

[71] Tomasz Burzykowski, Geert Molenberghs, Marc Buyse, Helena Geys, and Didier Renard. Validation of surrogate end points in multiple randomized clinical trials with failure time end points. Applied Statistics, 50(4):??, ???? 2001. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hazelton:2001:EOD

[72] Martin L. Hazelton. Estimation of origin–destination trip rates in Leicester. Applied Statistics, 50(4):??, ???? 2001. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Thomas:2001:MCC

[73] Hoben Thomas and Thomas P. Hettmansperger. Modelling change in cognitive understanding with finite mixtures. Applied Statistics, 50(4):??, ???? 2001. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lawless:2001:ARF

[74] J. F. Lawless, M. B. Wigg, S. Tuli, J. Drake, and M. Lamberti-Pasculli. Analysis of repeated failures or durations, with application to shunt failures for patients with paediatric hydrocephalus. Applied Statistics, 50(4): ??, ???? 2001. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Leong:2001:ICC

[75] Traci Leong, Stuart R. Lipsitz, and Joseph G. Ibrahim. Incomplete covariates in the Cox model with applications to biological marker data. Applied Statistics, 50(4):??, ???? 2001. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Polzehl:2001:FDM

[76] Jörg Polzehl and Vladimir G. Spokoiny. Functional and dynamic magnetic resonance imaging using vector adaptive weights smoothing. Applied Statistics, 50(4):??, ???? 2001. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Nicholls:2001:RDT

[77] Geoff Nicholls and Martin Jones. Radiocarbon dating with temporal order constraints. Applied Statistics, 50(4): ??, ???? 2001. CODEN APSTAG.

ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:2001:IAV

[78] Anonymous. Index of authors, volume 50, 2001. Applied Statistics, 50 (4):??, ???? 2001. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hogan:2002:HMA

[79] Joseph W. Hogan and Michael J. Daniels. A hierarchical modelling approach to analysing longitudinal data with drop-out and non-compliance, with application to an equivalence trial in paediatric acquired immune deficiency syndrome. Applied Statistics, 51(1): ??, ???? 2002. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Theobald:2002:BCC

[80] Chris M. Theobald and Mike Talbot. The Bayesian choice of crop variety and fertilizer dose. Applied Statistics, 51 (1):??, ???? 2002. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Huang:2002:BMM

[81] Guan-Hua Huang, Karen Bandeen-Roche, and Gary S. Rubin. Building marginal models for multiple ordinal measurements. Applied Statistics, 51 (1):??, ???? 2002. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Wang:2002:SRS

[82] Jixian Wang. Sample reuse simulation in optimal design for Tmax in pharmacokinetic experiments. Applied Statistics, 51 (1):??, ???? 2002. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Reboussin:2002:LCM

[83] Beth A. Reboussin, Michael E. Miller, Kurt K. Lohman, and Thomas R. Ten Have. Latent class models for longitudinal studies of the elderly with data missing at random. Applied Statistics, 51(1):??, ???? 2002. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kowalski:2002:GEE

[84] J. Kowalski and X. M. Tu. A generalized estimating equation approach to modelling incompatible data formats with covariate measurement error: application to human immunodeficiency virus immune markers. Applied Statistics, 51 (1):??, ???? 2002. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Korhonen:2002:EMR

[85] Pasi Korhonen and Juni Palmgren. Effect modification in a randomized trial under non-ignorable non-compliance: an application to the alpha-tocopherol beta-carotene study. Applied Statistics, 51(1):??, ???? 2002. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Chen:2002:BCR

[86] Ming-Hui Chen, David P. Harrington, and Joseph G. Ibrahim. Bayesian cure rate models for malignant melanoma: a case-study of Eastern Cooperative Oncology Group trial E1690. Applied Statistics, 51(2):135–150, ???? 2002.

CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bray:2002:AMC

[87] Isabelle Bray. Application of Markov chain Monte Carlo methods to projecting cancer incidence and mortality. Applied Statistics, 51(2):151–164, ???? 2002. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Foxall:2002:NMA

[88] Rob Foxall and Adrian Baddeley. Nonparametric measures of association between a spatial point process and a random set, with geological applications. Applied Statistics, 51(2):165–182, ???? 2002. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lucy:2002:NCA

[89] D. Lucy, R. G. Aykroyd, and A. M. Pollard. Nonparametric calibration for age estimation. Applied Statistics, 51(2): 183–196, ???? 2002. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Myles:2002:MCM

[90] Jonathan P. Myles, Stephen W. Duffy, Teresa C. Prevost, Nicholas E. Day, Matti Hakama, and Tiina Salminen. Markov chain Monte Carlo inference applied to a complex model for the longitudinal analysis of breast patterns. Applied Statistics, 51(2):197–207, ???? 2002. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Glasbey:2002:MPE

[91] C. A. Glasbey and M. J. Young. Maximum a posteriori estimation of image boundaries by dynamic programming.

Applied Statistics, 51(2):209–221, ???? 2002. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Arbia:2002:ASS

[92] Giuseppe Arbia and Giovanni Lafratta. Anisotropic spatial sampling designs for urban pollution. Applied Statistics, 51 (2):223–234, ???? 2002. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Duffy:2002:CNC

[93] Stephen W. Duffy, Jack Cuzick, Laszlo Tabar, Bedrich Vitak, Tony Hsiu-Hsi Chen, Ming-Fang Yen, and Robert A. Smith. Correcting for non-compliance bias in case-control studies to evaluate cancer screening programmes. Applied Statistics, 51(2):235–243, ???? 2002. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Wolfe:2002:MSU

[94] Rory Wolfe and David Firth. Modelling subjective use of an ordinal response scale in a many period crossover experiment. Applied Statistics, 51(2): 245–255, ???? 2002. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Fernandez:2002:BMC

[95] Carmen Fernández, Eduardo Ley, and Mark F. J. Steel. Bayesian modelling of catch in a north-west Atlantic fishery. Applied Statistics, 51(3):257–280, ???? 2002. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Horton:2002:MLE

[96] Nicholas J. Horton and Garrett M. Fitzmaurice. Maximum likelihood estimation of bivariate logistic models for incomplete responses with indicators of ignorable and non-ignorable missingness. Applied Statistics, 51(3):281–295, ???? 2002. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Wu:2002:MTD

[97] Lang Wu and Hulin Wu. Missing time-dependent covariates in human immunodeficiency virus dynamic models. Applied Statistics, 51(3):297–318, ???? 2002. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Francis:2002:APR

[98] Brian Francis, Regina Dittrich, Reinhold Hatzinger, and Roger Penn. Analysing partial ranks by using smoothed paired comparison methods: an investigation of value orientation in Europe. Applied Statistics, 51(3):319–336, ???? 2002. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Tatsuoka:2002:DAM

[99] Curtis Tatsuoka. Data analytic methods for latent partially ordered classification models. Applied Statistics, 51(3): 337–350, ???? 2002. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Shaddick:2002:MDM

[100] Gavin Shaddick and Jon Wakefield. Modelling daily multivariate pollutant data at multiple sites. Applied Statistics, 51(3):351–372, ???? 2002. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Aitchison:2002:BCD

[101] John Aitchison and Michael Greenacre.

Biplots of compositional data. Applied Statistics, 51(4):375–392, ???? 2002. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Baxter:2002:DRS

[102] Paul D. Baxter and Graham J. G.

Upton. Denoising radiocommunications signals by using iterative wavelet shrinkage. Applied Statistics, 51(4):393– 403, ???? 2002. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lambert:2002:MMA

[103] Philippe Lambert. A mixture model to assess the effect of hormonal stimulation on the development of follicles in prepubertal heifers. Applied Statistics, 51(4): 405–420, ???? 2002. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hemming:2002:PDS

[104] K. Hemming and J. E. H. Shaw. A parametric dynamic survival model applied to breast cancer survival times. Applied Statistics, 51(4):421–435, ???? 2002. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Robin:2002:CPM

[105] Stéphane Robin. A compound Poisson model for word occurrences in DNA sequences. Applied Statistics, 51(4):437– 451, ???? 2002. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Chiogna:2002:DGL

[106] Monica Chiogna and Carlo Gaetan. Dynamic generalized linear models with application to environmental epidemiology. Applied Statistics, 51(4):453–468, ???? 2002. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Cadigan:2002:GLI

[107] N. G. Cadigan and P. J. Farrell. Generalized local influence with applications to fish stock cohort analysis. Applied Statistics, 51(4):469–483, ???? 2002. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Gangnon:2002:MDE

[108] Ronald E. Gangnon and William N. King. Minimum distance estimation of the distribution functions of stochastically ordered random variables. Applied Statistics, 51(4):485–492, ???? 2002. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Diggle:2002:CMG

[109] Peter Diggle, Rana Moyeed, Barry Rowlingson, and Madeleine Thomson. Childhood malaria in the Gambia: A case-study in model-based geostatistics. Applied Statistics, 51(4):493–506, ???? 2002. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Paik:2002:C

[110] Myunghee Cho Paik. Correspondence. Applied Statistics, 51(4):507–508, ???? 2002. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:2002:IAV

[111] Anonymous. Index of authors, volume 51, 2002. Applied Statistics, 51(4):509– 511, ???? 2002. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kammann:2003:GM

[112] E. E. Kammann and M. P. Wand.

Geoadditive models. Applied Statistics, 52(1):1–18, January 2003. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Chen:2003:SOA

[113] Victoria C. P. Chen, Dirk Günther, and Ellis L. Johnson. Solving for an optimal airline yield management policy via statistical learning. Applied Statistics, 52 (1):19–30, January 2003. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Dryden:2003:BTS

[114] Ian L. Dryden, Mark R. Scarr, and Charles C. Taylor. Bayesian texture segmentation of weed and crop images using reversible jump Markov chain Monte Carlo methods. Applied Statistics, 52 (1):31–50, January 2003. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kuk:2003:GEE

[115] Anthony Y. C. Kuk. A generalized estimating equation approach to modelling foetal response in developmental toxicity studies when the number of implants is dose dependent. Applied Statistics, 52 (1):51–61, January 2003. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Georgiadis:2003:CAE

[116] Marios P. Georgiadis, Wesley O. Johnson, Ian A. Gardner, and Ramanpreet Singh. Correlation-adjusted estimation of sensitivity and specificity of two diagnostic tests. Applied Statistics, 52(1):63– 76, January 2003. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Basu:2003:BAC

[117] Sanjib Basu, Ananda Sen, and Mousumi Banerjee. Bayesian analysis of competing risks with partially masked cause of failure. Applied Statistics, 52(1):77– 93, January 2003. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Besbeas:2003:EIA

[118] P. Besbeas, J.-D. Lebreton, and B. J. T. Morgan. The efficient integration of abundance and demographic data. Applied Statistics, 52(1):95–102, January 2003. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Allen:2003:EDC

[119] Theodore T. Allen, Liyang Yu, and John Schmitz. An experimental design criterion for minimizing meta-model prediction errors applied to die casting process design. Applied Statistics, 52(1): 103–117, January 2003. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Renshaw:2003:LCM

[120] Arthur Renshaw and Steven Haberman. Lee–Carter mortality forecasting: a parallel generalized linear modelling approach for England and Wales mortality projections. Applied Statistics, 52(1):

119–137, January 2003. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Leung:2003:ESU

[121] Denis Heng-Yan Leung, You-Gan Wang, and David Amar. Early stopping by using stochastic curtailment in a three-arm sequential trial. Applied Statistics, 52 (2):139–152, May 2003. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kwong:2003:CPM

[122] G. P. S. Kwong and J. L. Hutton.

Choice of parametric models in survival analysis: applications to monotherapy for epilepsy and cerebral palsy. Applied Statistics, 52(2):153–168, May 2003. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Knorr-Held:2003:HMS

[123] Leonhard Knorr-Held and Sylvia Richardson. A hierarchical model for space-time surveillance data on meningococcal disease incidence. Applied Statistics, 52(2):169–183, May 2003. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Linardakis:2003:AAM

[124] Michalis Linardakis and Petros Dellaportas. Assessment of Athens’s metro passenger behaviour via a multiranked probit model. Applied Statistics, 52(2): 185–200, May 2003. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Sebastiani:2003:MRF

[125] Maria Rita Sebastiani. Markov randomfield models for estimating local labour markets. Applied Statistics, 52(2):201– 211, May 2003. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bockenholt:2003:ASD

[126] Ulf Böckenholt. Analysing state dependences in emotional experiences by dynamic count data models. Applied Statistics, 52(2):213–226, May 2003. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Upton:2003:RMC

[127] Graham J. G. Upton, Kenneth A. Hickey, and Aaron Stallard. Regression models for cyclic data. Applied Statistics, 52(2):227–235, May 2003. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Chau:2003:RIH

[128] P. H. Chau, Paul S. F. Yip, and Jisheng S. Cui. Reconstructing the incidence of human immunodeficiency virus (HIV) in Hong Kong by using data from HIV positive tests and diagnoses of acquired immune deficiency syndrome. Applied Statistics, 52(2):237–248, May 2003. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Sinha:2003:MAL

[129] Debajyoti Sinha, Kauhsik Patra, and Dipak K. Dey. Modelling accelerated life test data by using a Bayesian approach. Applied Statistics, 52(2):249–259, May 2003. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Atkinson:2003:HRT

[130] Anthony C. Atkinson. Horwitz’s Rule, transforming both sides and the design of experiments for mechanistic models. Applied Statistics, 52(3):261– 278, August 2003. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Battaglia:2003:FCI

[131] Francesco Battaglia and Livio Fenga.

Forecasting composite indicators with anticipated information: An application to the industrial production index. Applied Statistics, 52(3):279–290, August 2003. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Mezzetti:2003:BCM

[132] Maura Mezzetti, Joseph G. Ibrahim, Frédéric Y. Bois, Louise M. Ryan, Long Ngo, and Thomas J. Smith. A Bayesian compartmental model for the evaluation of 1,3-butadiene metabolism. Applied Statistics, 52(3):291–305, August 2003. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Chave:2003:BIR

[133] Alan D. Chave and David J. Thomson.

A bounded influence regression estimator based on the statistics of the hat matrix. Applied Statistics, 52(3):307– 322, August 2003. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hoggart:2003:BKD

[134] C. J. Hoggart, S. G. Walker, and A. F. M. Smith. Bivariate kurtotic distributions of garment fibre data. Applied Statistics, 52(3):323–335, August 2003. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Heffernan:2003:EVA

[135] Janet E. Heffernan and Jonathan A. Tawn. An extreme value analysis for the investigation into the sinking of the M. V. Derbyshire. Applied Statistics, 52 (3):337–354, August 2003. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Duchateau:2003:ERA

[136] Luc Duchateau, Paul Janssen, Iva Kezic, and Catherine Fortpied. Evolution of recurrent asthma event rate over time in frailty models. Applied Statistics, 52 (3):355–363, August 2003. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Gadbury:2003:RTS

[137] Gary L. UNKNOWN Gadbury, Grier P. Page, Moonseong Heo, John D. Mountz, and David B. Allison. Randomization tests for small samples: An application for genetic expression data. Applied Statistics, 52(3):365–376, August 2003. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hirst:2003:HMA

[138] David Hirst, Geir Storvik, and Anne Randi Syversveen. A hierarchical modelling approach to combining environmental data at different scales. Applied Statistics, 52(3):377–390, August 2003. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lin:2003:CPC

[139] Hung-Mo Lin, John M. Williamson, and Stuart R. Lipsitz. Calculating power for the comparison of dependent κcoefficients. Applied Statistics, 52(4):

??, November 2003. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Coles:2003:ACT

[140] Stuart Coles and Luis Pericchi. Anticipating catastrophes through extreme value modelling. Applied Statistics, 52 (4):??, November 2003. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Brockwell:2003:CMA

[141] A. E. Brockwell, N. H. Chan, and P. K. Lee. A class of models for aggregated traffic volume time series. Applied Statistics, 52(4):??, November 2003. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Carpenter:2003:NBP

[142] James R. Carpenter, Harvey Goldstein, and Jon Rasbash. A novel bootstrap procedure for assessing the relationship between class size and achievement. Applied Statistics, 52(4):??, November 2003. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Farewell:2003:TSA

[143] Vernon T. Farewell, Jerald F. Lawless, Dafna D. Gladman, and Murray B. Urowitz. Tracing studies and analysis of the effect of loss to follow-up on mortality estimation from patient registry data. Applied Statistics, 52(4): ??, November 2003. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Berger:2003:VEL

[144] Yves G. Berger and Chris J. Skinner.

Variance estimation for a low income proportion. Applied Statistics, 52(4): ??, November 2003. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Shkedy:2003:MFI

[145] Ziv Shkedy, Marc Aerts, Geert Molenberghs, Philippe Beutels, and Pierre Van Damme. Modelling forces of infection by using monotone local polynomials. Applied Statistics, 52(4):??, November 2003. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Allcroft:2003:LGM

[146] David J. Allcroft and Chris A. Glasbey. A latent Gaussian Markov randomfield model for spatiotemporal rainfall disaggregation. Applied Statistics, 52(4): ??, November 2003. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ogata:2003:MHS

[147] Yosihiko Ogata, Koichi Katsura, and Masaharu Tanemura. Modelling heterogeneous space–time occurrences of earthquakes and its residual analysis. Applied Statistics, 52(4):??, November 2003. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:2003:IAV

[148] Anonymous. Index of authors, volume 52, 2003. Applied Statistics, 52(4):??, November 2003. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hirst:2004:ECA

[149] David Hirst, Sondre Aanes, Geir Storvik, Ragnar Bang Huseby, and Ingunn Fride Tvete. Estimating catch at age from market sampling data by using a Bayesian hierarchical model. Applied Statistics, 53(1):1–14, February 2004. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Oh:2004:PAV

[150] Hee-Seok Oh, Doug Nychka, Tim Brown, and Paul Charbonneau. Period analysis of variable stars by robust smoothing. Applied Statistics, 53(1):15– 30, February 2004. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Al-Awadhi:2004:SIA

[151] Fahimah Al-Awadhi, Christopher Jennison, and Merrilee Hurn. Statistical image analysis for a confocal microscopy two-dimensional section of cartilage growth. Applied Statistics, 53(1): 31–49, February 2004. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Gasparini:2004:RSI

[152] Mauro Gasparini, Harald Nusser, and Jeffrey Eisele. Repeated screening with inspection error and no false positive results with application to pharmaceutical pill production. Applied Statistics, 53 (1):51–62, February 2004. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Landau:2004:TDT

[153] S. Landau, I. C. Ellison-Wright, and E. T. Bullmore. Tests for a difference in timing of physiological response between two brain regions measured by using functional magnetic resonance imaging. Applied Statistics, 53(1):63– 82, February 2004. CODEN APSTAG.

ISSN 0035-9254 (print), 1467-9876 (electronic).

Oakley:2004:EPU

[154] Jeremy Oakley. Estimating percentiles of uncertain computer code outputs. Applied Statistics, 53(1):83–93, February 2004. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Beasley:2004:CIN

[155] T. Mark Beasley, Grier P. Page, Jaap P. L. Brand, Gary L. Gadbury, John D. Mountz, and David B. Allison. Chebyshev’s inequality for nonparametric testing with small N and α in microarray research. Applied Statistics, 53(1):95– 108, February 2004. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Aitken:2004:ETE

[156] C. G. G. Aitken and D. Lucy. Evaluation of trace evidence in the form of multivariate data. Applied Statistics, 53(1): 109–122, February 2004. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Catchpole:2004:CMI

[157] E. A. Catchpole, B. J. T. Morgan, and T. Coulson. Conditional methodology for individual case history data. Applied Statistics, 53(1):123–131, February 2004. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Godolphin:2004:SPP

[158] J. D. Godolphin. Simple pilot procedures for the avoidance of disconnected experimental designs. Applied Statistics, 53(1):133–147, February 2004. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Jackson:2004:MLD

[159] Christopher H. Jackson and Linda D.

Sharples. Models for longitudinal data with censored changepoints. Applied Statistics, 53(1):149–162, February 2004. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Theobald:2004:PCA

[160] C. M. Theobald, C. A. Glasbey, G. W.

Horgan, and C. D. Robinson. Principal component analysis of landmarks from reversible images. Applied Statistics, 53(1):163–175, February 2004. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

OBrien:2004:ALM

[161] Liam M. O’Brien and Garrett M.

Fitzmaurice. Analysis of longitudinal multiple-source binary data using generalized estimating equations. Applied Statistics, 53(1):177–193, February 2004. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Stephens:2004:BAD

[162] David Stephens and Martin Crowder.

Bayesian analysis of discrete time warranty data. Applied Statistics, 53(1): 195–217, February 2004. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Mateu:2004:UDL

[163] Jorge Mateu, Francisco Montes, and Mario Plaza. The 1970 US draft lottery revisited: a spatial analysis. Applied Statistics, 53(1):219–229, February 2004. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Huerta:2004:SMM

[164] Gabriel Huerta, Bruno Sansó, and Jonathan R. Stroud. A spatiotemporal model for Mexico City ozone levels. Applied Statistics, 53(2):231–248, April 2004. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Gill:2004:BAD

[165] Paramjit S. Gill and Tim B. Swartz. Bayesian analysis of directed graphs data with applications to social networks. Applied Statistics, 53(2):249–260, April 2004.

Brazzale:2004:HMA

[166] Alessandra Rosalba Brazzale, Alberto Salvan, and Stefano Roletti. A hierarchical modelling approach for measuring reliability of and agreement between two types of magnetic field dosimeter. Applied Statistics, 53(2):261–278, April 2004.

Lipsitz:2004:GLM

[167] Stuart Lipsitz, Michael Parzen, Sundar Natarajan, Joseph Ibrahim, and Garrett Fitzmaurice. Generalized linear models with a coarsened covariate. Applied Statistics, 53(2):279–292, April 2004.

Herring:2004:NIM

[168] Amy H. Herring, Joseph G. Ibrahim, and Stuart R. Lipsitz. Non-ignorable missing covariate data in survival analysis: a case-study of an International Breast Cancer Study Group trial. Applied Statistics, 53(2):293–310, April 2004.

Foulkes:2004:CGG

[169] A. S. Foulkes, V. De Gruttola, and K. Hertogs. Combining genotype groups and recursive partitioning: an application to human immunodeficiency virus type 1 genetics data. Applied Statistics, 53(2):311–323, April 2004.

Bingham:2004:DFF

[170] D. R. Bingham, E. D. Schoen, and R. R.

Sitter. Designing fractional factorial split-plot experiments with few wholeplot factors. Applied Statistics, 53(2): 325–339, April 2004.

Rutter:2004:BME

[171] Carolyn M. Rutter and Gregory Simon.

A Bayesian method for estimating the accuracy of recalled depression. Applied Statistics, 53(2):341–353, April 2004.

Kauermann:2004:TPN

[172] Göran Kauermann and Renate Ortlieb.

Temporal pattern in number of staff on sick leave: the effect of downsizing. Applied Statistics, 53(2):355–367, April 2004.

Kuk:2004:LBA

[173] Anthony Y. C. Kuk. A litter-based approach to risk assessment in developmental toxicity studies via a power family of completely monotone functions. Applied Statistics, 53(2):369–386, April 2004.

Zhang:2004:SMS

[174] Wenyang Zhang and Fiona Steele.

A semiparametric multilevel survival model. Applied Statistics, 53(2):387– 404, April 2004.

Caussinus:2004:DCA

[175] Henri Caussinus and Olivier Mestre. Detection and correction of artificial shifts in climate series. Applied Statistics, 53 (3):405–425, August 2004. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Santis:2004:OPS

[176] Fulvio De Santis, Marco Perone Pacifico, and Valeria Sambucini. Optimal predictive sample size for case-control studies. Applied Statistics, 53(3):427–441, August 2004.

Chakraborti:2004:CDF

[177] S. Chakraborti, P. van der Laan, and M. A. van de Wiel. A class of distribution-free control charts. Applied Statistics, 53(3):443–462, August 2004.

Cruz-Medina:2004:SMM

[178] I. R. Cruz-Medina, T. P. Hettmansperger, and H. Thomas. Semiparametric mixture models and repeated measures: the multinomial cut point model. Applied Statistics, 53(3):463–474, August 2004.

dePasquale:2004:BAD

[179] Francesco de Pasquale, Piero Barone, Giovanni Sebastiani, and Julian Stander. Bayesian analysis of dynamic magnetic resonance breast images. Applied Statistics, 53(3):475–493, August 2004.

Nummi:2004:EPL

[180] Tapio Nummi and Jyrki Möttönen. Estimation and prediction for low degree polynomial models under measurement errors with an application to forest harvesters. Applied Statistics, 53(3):495– 505, August 2004.

Scott:2004:MGD

[181] Marc A. Scott, Robert G. Norman, and Kenneth I. Berger. Modelling growth and decline in lung function in Duchenne’s muscular dystrophy with an augmented linear mixed effects model. Applied Statistics, 53(3):507–521, August 2004.

Lindsey:2004:ALO

[182] P. J. Lindsey and J. Kaufmann. Analysis of a longitudinal ordinal response clinical trial using dynamic models. Applied Statistics, 53(3):523–537, August 2004.

Rousson:2004:SCA

[183] Valentin Rousson and Theo Gasser.

Simple component analysis. Applied Statistics, 53(4):539–555, November 2004. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Muller:2004:TCS

[184] T. G. Müller, D. Faller, J. Timmer, I. Swameye, O. Sandra, and U. Klingmüller. Tests for cycling in a signalling pathway. Applied Statistics, 53(4):557–568, November 2004.

Berger:2004:RDL

[185] Martijn P. F. Berger and Frans E. S.

Tan. Robust designs for linear mixed effects models. Applied Statistics, 53(4): 569–581, November 2004.

Politis:2004:BUA

[186] Konstadinos Politis and Lennart Robertson. Bayesian updating of atmospheric dispersion after a nuclear accident. Applied Statistics, 53(4):583–600, November 2004.

Harkanen:2004:TIP

[187] T. Härkänen and E. Arjas. Tumour incidence, prevalence and lethality estimation in the absence of cause-of-death information. Applied Statistics, 53(4):601– 617, November 2004.

Lloyd:2004:RBE

[188] Chris J. Lloyd and Donald J. Frommer. Regression-based estimation of the false negative fraction when multiple negatives are unverified. Applied Statistics, 53(4):619–631, November 2004.

Berzuini:2004:EPA

[189] Carlo Berzuini and Claudia Allemani. Effectiveness of potent antiretroviral therapy on progression of human immunodeficiency virus: Bayesian modelling and model checking via counterfactual replicates. Applied Statistics, 53 (4):633–650, November 2004.

Antognini:2004:NBC

[190] Alessandro Baldi Antognini and Alessandra Giovagnoli. A new ‘biased coin design’ for the sequential allocation of two treatments. Applied Statistics, 53(4): 651–664, November 2004.

Aitken:2004:CET

[191] C. G. G. Aitken and D. Lucy. Corrigendum: Evaluation of trace evidence in the form of multivariate data. Applied Statistics, 53(4):665–666, November 2004.

Anonymous:2004:IAV

[192] Anonymous. Index of authors, volume 53, 2004. Applied Statistics, 53(4):667– 669, November 2004.

Gelfand:2005:MSD

[193] Alan E. Gelfand, Alexandra M. Schmidt, Shanshan Wu, John A. Silander, Jr., Andrew Latimer, and Anthony G. Rebelo. Modelling species diversity through species level hierarchical modelling. Applied Statistics, 54(1):1–20, January 2005. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Huang:2005:AIE

[194] Li-Shan Huang, Hongkun Wang, and Christopher Cox. Assessing interaction effects in linear measurement error models. Applied Statistics, 54(1):21–30, January 2005.

Delmar:2005:MMV

[195] Paul Delmar, Stéphane Robin, Diana Tronik-Le Roux, and Jean Jacques Daudin. Mixture model on the variance for the differential analysis of gene expression data. Applied Statistics, 54(1): 31–50, January 2005.

Peddada:2005:SAQ

[196] Shyamal D. Peddada, Gregg E. Dinse, and Joseph K. Haseman. A survivaladjusted quantal response test for comparing tumour incidence rates. Applied Statistics, 54(1):51–61, January 2005.

Cox:2005:LQL

[197] Christopher Cox. Limits of quantitation for laboratory assays. Applied Statistics, 54(1):63–76, January 2005.

Mwalili:2005:BOL

[198] Samuel M. Mwalili, Emmanuel Lesaffre, and Dominique Declerck. A Bayesian ordinal logistic regression model to correct for interobserver measurement error in a geographical oral health study. Applied Statistics, 54(1):77–93, January 2005.

Denteneer:2005:NEC

[199] D. Denteneer and M. S. Keane. Nonparametric estimation of a change in defect intensity. Applied Statistics, 54(1): 95–114, January 2005.

Stare:2005:IMR

[200] Janez Stare, Robin Henderson, and Maja Pohar. An individual measure of relative survival. Applied Statistics, 54 (1):115–126, January 2005.

Shmueli:2005:UDF

[201] Galit Shmueli, Thomas P. Minka, Joseph B. Kadane, Sharad Borle, and Peter Boatwright. A useful distribution for fitting discrete data: Revival of the Conway–Maxwell–Poisson distribution. Applied Statistics, 54(1):127– 142, January 2005. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://www.jstor.org/
stable/3592603.

Gilbert:2005:MFD

[202] Peter B. Gilbert. A modified false discovery rate multiple-comparisons procedure for discrete data, applied to human immunodeficiency virus genetics. Applied Statistics, 54(1):143–158, January 2005.

Peters:2005:BMC

[203] Jaime L. Peters, Lesley Rushton, Alex J. Sutton, David R. Jones, Keith R. Abrams, and Moira A. Mugglestone. Bayesian methods for the cross-design synthesis of epidemiological and toxicological evidence. Applied Statistics, 54 (1):159–172, January 2005.

Alonzo:2005:AAC

[204] Todd A. Alonzo and Margaret Sullivan Pepe. Assessing accuracy of a continuous screening test in the presence of verification bias. Applied Statistics, 54(1): 173–190, January 2005.

Verzilli:2005:HBM

[205] Claudio J. Verzilli, John C. Whittaker, Nigel Stallard, and Daniel Chasman. A hierarchical Bayesian model for predicting the functional consequences of amino-acid polymorphisms. Applied Statistics, 54(1):191–206, January 2005.

Chavez-Demoulin:2005:GAM

[206] V. Chavez-Demoulin and A. C. Davison.

Generalized additive modelling of sample extremes. Applied Statistics, 54(1): 207–222, January 2005.

Sahu:2005:BKK

[207] Sujit K. Sahu and Kanti V. Mardia.

A Bayesian kriged Kalman model for short-term forecasting of air pollution levels. Applied Statistics, 54(1):223–244, January 2005.

Piegorsch:2005:LDR

[208] Walter W. Piegorsch, R. Webster West, Wei Pan, and Ralph L. Kodell. Low dose risk estimation via simultaneous statistical inferences. Applied Statistics, 54(1): 245–258, January 2005.

Wannemuehler:2005:UMC

[209] Kathleen A. Wannemuehler and Robert H.

Lyles. A unified model for covariate measurement error adjustment in an occupational health study while accounting for non-detectable exposures. Applied Statistics, 54(1):259–271, January

Saebo:2005:DRM

[210] S. Sæbø, T. Almøy, and A. H. Aastveit. Disease resistance modelled as firstpassage times of genetically dependent stochastic processes. Applied Statistics, 54(1):273–285, January 2005.

Rambharat:2005:TAM

[211] B. Ricky Rambharat, Anthony E. Brockwell, and Duane J. Seppi. A threshold autoregressive model for wholesale electricity prices. Applied Statistics, 54(2):287–299, April 2005. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bowman:2005:JML

[212] F. DuBois Bowman and Amita K. Manatunga. A joint model for longitudinal data profiles and associated event risks with application to a depression study. Applied Statistics, 54(2):301–316, April 2005.

Martin-Magniette:2005:NEH

[213] M. L. Martin-Magniette. Nonparametric estimation of the hazard function by using a model selection method: estimation of cancer deaths in Hiroshima atomic bomb survivors. Applied Statistics, 54(2):317–331, April 2005.

DeRidder:2005:ESO

[214] Joris De Ridder, Geert Molenberghs, and Conny Aerts. Estimating stellaroscillation-related parameters and their uncertainties with the moment method. Applied Statistics, 54(2):333–348, April 2005.

Hohle:2005:IDT

[215] Michael Höhle, Erik Jørgensen, and Philip D. O’Neill. Inference in disease transmission experiments by using stochastic epidemic models. Applied Statistics, 54(2):349–366, April 2005.

Sidik:2005:SHV

[216] Kurex Sidik and Jeffrey N. Jonkman.

Simple heterogeneity variance estimation for meta-analysis. Applied Statistics, 54(2):367–384, April 2005.

Welton:2005:MTI

[217] N. J. Welton and A. E. Ades. A model of toxoplasmosis incidence in the UK: evidence synthesis and consistency of evidence. Applied Statistics, 54(2):385–404, April 2005.

Pantazis:2005:BML

[218] N. Pantazis, G. Touloumi, A. S. Walker, and A. G. Babiker. Bivariate modelling of longitudinal measurements of two human immunodeficiency type 1 disease progression markers in the presence of informative drop-outs. Applied Statistics, 54(2):405–423, April 2005.

Wilson:2005:HMO

[219] Simon P. Wilson. Hierarchical modelling of orthopaedic hip replacement damage accumulation and reliability. Applied Statistics, 54(2):425–441, April 2005.

Huang:2005:DAV

[220] Deng Huang and Theodore T. Allen. Design and analysis of variable fidelity experimentation applied to engine valve heat treatment process design. Applied Statistics, 54(2):443–463, April 2005.

Tatsuoka:2005:CDA

[221] Curtis Tatsuoka. Corrigendum: Data analytic methods for latent partially ordered classification models. Applied Statistics, 54(2):465–467, April 2005.

Kou:2005:BAS

[222] S. C. Kou, X. Sunney Xie, and Jun S. Liu. Bayesian analysis of single-molecule experimental data. Applied Statistics, 54 (3):469–506, June 2005. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Rigby:2005:GAM

[223] R. A. Rigby and D. M. Stasinopoulos. Generalized additive models for location, scale and shape. Applied Statistics, 54(3):507–554, June 2005.

Salim:2005:MAB

[224] A. Salim, Y. Pawitan, and K. Bond. Modelling association between two irregularly observed spatiotemporal processes by using maximum covariance analysis. Applied Statistics, 54(3):555– 573, June 2005.

Morton:2005:DTM

[225] Alexander Morton and Bärbel F. Finkenstädt. Discrete time modelling of disease incidence time series by using Markov chain Monte Carlo methods. Applied Statistics, 54(3):575–594, June 2005.

Elliott:2005:OCR

[226] Michael R. Elliott and William W. Davis. Obtaining cancer risk factor prevalence estimates in small areas: combining data from two surveys. Applied Statistics, 54(3):595–609, June 2005.

Muller:2005:NBM

[227] Peter Müller, Gary L. Rosner, Maria De Iorio, and Steven MacEachern. A nonparametric Bayesian model for inference in related longitudinal studies. Applied Statistics, 54(3):611–626, June 2005.

Do:2005:BMM

[228] Kim-Anh Do, Peter Müller, and Feng Tang. A Bayesian mixture model for differential gene expression. Applied Statistics, 54(3):627–644, June 2005.

Diggle:2005:NES

[229] Peter Diggle, Pingping Zheng, and Peter Durr. Nonparametric estimation of spatial segregation in a multivariate point process: bovine tuberculosis in Cornwall, UK. Applied Statistics, 54(3):645– 658, June 2005.

Xia:2005:SED

[230] Yingcun Xia, Julia R. Gog, and Bryan T. Grenfell. Semiparametric estimation of the duration of immunity from infectious disease time series: influenza as a case-study. Applied Statistics, 54 (3):659–672, June 2005.

Dryden:2005:NSS

[231] I. L. Dryden, L. Márkus, C. C. Taylor, and J. Kovács. Non-stationary spatiotemporal analysis of karst water levels. Applied Statistics, 54(3):673–690, June 2005.

Agresti:2005:MTC

[232] Alan Agresti and Bernhard Klingenberg.

Multivariate tests comparing binomial probabilities, with application to safety studies for drugs. Applied Statistics, 54 (4):691–706, August 2005. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Albert:2005:MLS

[233] Paul S. Albert and Joannie Shen. Modelling longitudinal semicontinuous emesis volume data with serial correlation in an acupuncture clinical trial. Applied Statistics, 54(4):707–720, August 2005.

Bohning:2005:NML

[234] Dankmar Böhning and Dieter Schön. Nonparametric maximum likelihood estimation of population size based on the counting distribution. Applied Statistics, 54(4):721–737, August 2005.

Alioum:2005:MAE

[235] Ahmadou Alioum, Daniel Commenges, Rodolphe Thiébaut, and François Dabis. A multistate approach for estimating the incidence of human immunodeficiency virus by using data from a prevalent cohort study. Applied Statistics, 54(4): 739–752, August 2005.

Baxter:2005:DNR

[236] Paul D. Baxter and Graham J. G. Upton. Differentiating noisy radiocommunications signals: wavelet estimation of a derivative in the presence of heteroscedastic noise. Applied Statistics, 54 (4):753–767, August 2005.

Houseman:2005:RRM

[237] E. Andrés Houseman. A robust regression model for a first-order autoregressive time series with unequal spacing: application to water monitoring. Applied Statistics, 54(4):769–780, August 2005.

Meulders:2005:PFA

[238] Michel Meulders, Paul De Boeck, Iven Van Mechelen, and Andrew Gelman. Probabilistic feature analysis of facial perception of emotions. Applied Statistics, 54(4):781–793, August 2005.

Martel-Escobar:2005:AIH

[239] M. Martel-Escobar, F. J. Vázquez-Polo, and A. Hernández-Bastida. Analysing the independence hypothesis in models for rare errors: an application to auditing. Applied Statistics, 54(4):795–804, August 2005.

Blaauw:2005:RPC

[240] Maarten Blaauw and J. Andrés Christen. Radiocarbon peat chronologies and environmental change. Applied Statistics, 54(4):805–816, August 2005.

Wit:2005:NOD

[241] Ernst Wit, Agostino Nobile, and Raya Khanin. Near-optimal designs for dual channel microarray studies. Applied Statistics, 54(5):817–830, November 2005. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Chu:2005:CTC

[242] Haitao Chu, Lawrence H. Moulton, Wendy J. Mack, Douglas J. Passaro, Paulo F. Barroso, and Alvaro Muñoz. Correlating two continuous variables subject to detection limits in the context of mixture distributions. Applied Statistics, 54(5):831–845, November 2005.

Matthews:2005:AFA

[243] Abigail G. Matthews, Dianne M. Finkelstein, and Rebecca A. Betensky. Analysis of familial aggregation in the presence of varying family sizes. Applied Statistics, 54(5):847–862, November 2005.

Whittaker:2005:NTQ

[244] Joe Whittaker, Chris Whitehead, and Mark Somers. The neglog transformation and quantile regression for the analysis of a large credit scoring database. Applied Statistics, 54(5):863– 878, November 2005.

Kharroubi:2005:EUI

[245] Samer A. Kharroubi, Anthony O’Hagan, and John E. Brazier. Estimating utilities from individual health preference data: a nonparametric Bayesian method. Applied Statistics, 54(5):879–895, November 2005.

Wilson:2005:PFD

[246] Simon P. Wilson and Mark J. Costello. Predicting future discoveries of European marine species by using a nonhomogeneous renewal process. Applied Statistics, 54(5):897–918, November 2005.

Tiwari:2005:BMS

[247] Ram C. Tiwari, Kathleen A. Cronin, William Davis, Eric J. Feuer, Binbing Yu, and Siddhartha Chib. Bayesian model selection for join point regression with application to age-adjusted cancer rates. Applied Statistics, 54(5):919–939, November 2005.

Baker:2005:SWB

[248] Stuart G. Baker and Barnett S. Kramer. Statistics for weighing benefits and harms in a proposed genetic substudy of a randomized cancer prevention trial. Applied Statistics, 54(5):941–954, November 2005.

Bingham:2005:CDF

[249] D. R. Bingham, E. D. Schoen, and R. R. Sitter. Corrigendum: Designing fractional factorial split-plot experiments with few whole-plot factors. Applied Statistics, 54(5):955–958, November 2005.

Elliott:2005:COC

[250] Michael R. Elliott and William W.

Davis. Corrigendum: Obtaining cancer risk factor prevalence estimates in small areas: combining data from two surveys. Applied Statistics, 54(5):958, November 2005.

Anonymous:2005:IAV

[251] Anonymous. Index of authors, volume 54, 2005. Applied Statistics, 54(5):959– 962, November 2005.

Anonymous:2005:CV

[252] Anonymous. Contents of volume 54, 2005. Applied Statistics, 54(5):963–965, November 2005.

Dean:2006:UUD

[253] Nema Dean, Thomas Brendan Murphy, and Gerard Downey. Using unlabelled data to update classification rules with applications in food authenticity studies. Applied Statistics, 55(1):1–14, January 2006. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Stallard:2006:ENI

[254] Nigel Stallard, Mike B. Gravenor, and Robert N. Curnow. Estimating numbers of infectious units from serial dilution assays. Applied Statistics, 55(1):15–30, January 2006.

Cowling:2006:JME

[255] B. J. Cowling, J. L. Hutton, and J. E. H.

Shaw. Joint modelling of event counts and survival times. Applied Statistics, 55(1):31–39, January 2006.

Demyanov:2006:IEI

[256] V. Demyanov, S. N. Wood, and T. J. Kedwards. Improving ecological impact assessment by statistical data synthesis using process-based models. Applied Statistics, 55(1):41–62, January 2006.

Ridout:2006:MER

[257] Martin S. Ridout, Malcolm J. Faddy, and Michael G. Solomon. Modelling the effects of repellent chemicals on foraging bees. Applied Statistics, 55(1):63–75, January 2006.

Bock:2006:MAA

[258] Mitchum T. Bock and Adrian W. Bowman. On the measurement and analysis of asymmetry with applications to facial modelling. Applied Statistics, 55(1):77– 91, January 2006.

Verbeke:2006:CPC

[259] G. Verbeke, S. Fieuws, E. Lesaffre, B. S. Kato, M. D. Foreman, P. L. O. Broos, and K. Milisen. A comparison of procedures to correct for base-line differences in the analysis of continuous longitudinal data: a case-study. Applied Statistics, 55(1):93–101, January 2006.

Dryden:2006:ESA

[260] Ian L. Dryden and András Zempléni. Extreme shape analysis. Applied Statistics, 55(1):103–121, January 2006.

Newton:2006:APV

[261] Michael A. Newton and David I. Hastie. Assessing Poisson variation of intestinal tumour multiplicity in mice carrying a Robertsonian translocation. Applied Statistics, 55(1):123–138, January 2006.

Lee:2006:DHG

[262] Youngjo Lee and John A. Nelder. Double hierarchical generalized linear models (with discussion). Applied Statistics, 55(2):139–185, April 2006. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Jeong:2006:DPI

[263] Jong-Hyeon Jeong and Jason Fine. Direct parametric inference for the cumulative incidence function. Applied Statistics, 55(2):187–200, April 2006.

So:2006:BSS

[264] Mike K. P. So, Cathy W. S. Chen, and Feng-Chi Liu. Best subset selection of autoregressive models with exogenous variables and generalized autoregressive conditional heteroscedasticity errors. Applied Statistics, 55(2):201– 224, April 2006.

Baker:2006:POA

[265] Rose Baker and Philip Scarf. Predicting the outcomes of annual sporting contests. Applied Statistics, 55(2):225–239, April 2006.

Brooks:2006:DIP

[266] Chris Brooks and Melvin J. Hinich. Detecting intraday periodicities with application to high frequency exchange rates. Applied Statistics, 55(2):241–259, April 2006.

Dominici:2006:EPS

[267] Francesca Dominici, Scott L. Zeger, Giovanni Parmigiani, Joanne Katz, and Parul Christian. Estimating percentilespecific treatment effects in counterfactual models: a case-study of micronutrient supplementation, birth weight and infant mortality. Applied Statistics, 55 (2):261–280, April 2006.

Proietti:2006:DFA

[268] Tommaso Proietti and Filippo Moauro. Dynamic factor analysis with non-linear temporal aggregation constraints. Applied Statistics, 55(2):281–300, April 2006.

Pan:2006:IGF

[269] Wei Pan. Incorporating gene functional annotations in detecting differential gene expression. Applied Statistics, 55(3):301–316, May 2006. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Yang:2006:DEP

[270] Yang Yang, Ira M. Longini, Jr., and M. Elizabeth Halloran. Design and evaluation of prophylactic interventions using infectious disease incidence data from close contact groups. Applied Statistics, 55(3):317–330, May 2006.

Clarke:2006:CRS

[271] E. D. Clarke, D. C. Speirs, M. R. Heath, S. N. Wood, W. S. C. Gurney, and S. J. Holmes. Calibrating remotely sensed chlorophyll-a data by using penalized regression splines. Applied Statistics, 55 (3):331–353, May 2006. See corrigendum [285].

Arndt:2006:ESS

[272] Channing Arndt, Julia Kozlitina, and Paul V. Preckel. Efficient survey sampling of households via Gaussian quadrature. Applied Statistics, 55(3): 355–364, May 2006. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Demidenko:2006:ATR

[273] Eugene Demidenko. The assessment of tumour response to treatment. Applied Statistics, 55(3):365–377, May 2006.

Leung:2006:ASD

[274] Denis H. Y. Leung and Jing Qin.

Analysing survey data with incomplete responses by using a method based on empirical likelihood. Applied Statistics, 55(3):379–396, May 2006.

Belluzzo:2006:TST

[275] Walter Belluzzo. A two-sample test for threshold crossing latent variables with application to valuation studies. Applied Statistics, 55(3):397–406, May 2006.

Guillot:2006:IHS

[276] Gilles Guillot, Denis Kan-King-Yu, Joël Michelin, and Philippe Huet. Inference of a hidden spatial tessellation from multivariate data: application to the delineation of homogeneous regions in an agricultural field. Applied Statistics, 55 (3):407–430, May 2006.

Cadarso-Suarez:2006:FMN

[277] Carmen Cadarso-Suárez, Javier Roca-Pardiñas, Geert Molenberghs, Christel Faes, Verónica Nácher, Sabiela Ojeda, and Carlos Acuña. Flexible modelling of neuron firing rates across different experimental conditions: an application to neural activity in the prefrontal cortex during a discrimination task. Applied Statistics, 55(4):431–447, August 2006. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Fieuws:2006:HDM

[278] Steffen Fieuws, Geert Verbeke, Filip Boen, and Christophe Delecluse. High dimensional multivariate mixed models for binary questionnaire data. Applied Statistics, 55(4):449–460, August 2006.

Einbeck:2006:MBR

[279] Jochen Einbeck and Gerhard Tutz. Modelling beyond regression functions: an application of multimodal regression to speed-flow data. Applied Statistics, 55 (4):461–475, August 2006.

Cai:2006:AFS

[280] Liming Cai, Nathaniel Schenker, and James Lubitz. Analysis of functional status transitions by using a semi-Markov process model in the presence of leftcensored spells. Applied Statistics, 55(4): 477–491, August 2006.

Peddada:2006:TST

[281] Shyamal D. Peddada, Joseph K. Haseman, Xiaofeng Tan, and Greg Travlos. Tests for a simple tree order restriction with application to dose-response studies. Applied Statistics, 55(4):493–506, August 2006.

Parsons:2006:GEE

[282] N. R. Parsons, R. N. Edmondson, and S. G. Gilmour. A generalized estimating equation method for fitting autocorrelated ordinal score data with an application in horticultural research. Applied Statistics, 55(4):507–524, August 2006.

Spiess:2006:ETE

[283] Martin Spiess. Estimation of a twoequation panel model with mixed continuous and ordered categorical outcomes and missing data. Applied Statistics, 55(4):525–538, August 2006.

Yuan:2006:MRM

[284] Ao Yuan and George E. Bonney. A multivariate regression model for continuous genetic traits. Applied Statistics, 55 (4):539–550, August 2006. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Clarke:2006:CCR

[285] E. D. Clarke, D. C. Speirs, M. R. Heath, S. N. Wood, W. S. C. Gurney, and S. J. Holmes. Corrigendum: Calibrating remotely sensed chlorophyll-a data by using penalized regression splines. Applied Statistics, 55(4):551–552, August 2006. See [271].

Farrington:2006:SAC

[286] C. P. Farrington and H. J. Whitaker.

Semiparametric analysis of case series data. Applied Statistics, 55(5):553–594, November 2006. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Faes:2006:EHS

[287] Christel Faes, Niel Hens, Marc Aerts, Ziv Shkedy, Helena Geys, Koen Mintiens, Hans Laevens, and Frank Boelaert. Estimating herd-specific force of infection by using random-effects models for clustered binary data and monotone fractional polynomials. Applied Statistics, 55(5):595–613, November 2006.

Kowalski:2006:MCC

[288] Scott M. Kowalski, G. Geoffrey Vining, Douglas C. Montgomery, and Connie M. Borror. Modifying a central composite design to model the process mean and variance when there are hard-to-change factors. Applied Statistics, 55(5):615– 630, November 2006.

Fawcett:2006:HME

[289] Lee Fawcett and David Walshaw. A hierarchical model for extreme wind speeds. Applied Statistics, 55(5):631– 646, November 2006.

Goegebeur:2006:LIB

[290] Yuri Goegebeur, Paul De Boeck, Geert Molenberghs, and Guido del Pino. A local-influence-based diagnostic approach to a speeded item response theory model. Applied Statistics, 55(5):647– 676, November 2006.

Park:2006:PAM

[291] Yousung Park, Jai Won Choi, and Dong-Hee Lee. A parametric approach for measuring the effect of the 10th revision of the international classification of diseases. Applied Statistics, 55(5):677–697, November 2006.

Chen:2006:PDE

[292] Tao Chen, Julian Morris, and Elaine Martin. Probability density estimation via an infinite Gaussian mixture model: application to statistical process monitoring. Applied Statistics, 55(5):699– 715, November 2006.

Anonymous:2006:IAV

[293] Anonymous. Index of authors, volume 55, 2006. Applied Statistics, 55(5):717– 720, November 2006.

Jank:2007:MCE

[294] Wolfgang Jank and Galit Shmueli. Modelling concurrency of events in on-line auctions via spatiotemporal semiparametric models. Applied Statistics, 56(1): 1–27, January 2007. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Augustin:2007:SMN

[295] Nicole H. Augustin, Stefan Lang, Monica Musio, and Klaus Von Wilpert. A spatial model for the needle losses of pine-trees in the forests of Baden-Württemberg: an application of Bayesian structured additive regression. Applied Statistics, 56(1):29–50, January 2007.

Rodriguez-Avi:2007:GBB

[296] J. Rodríguez-Avi, A. Conde-Sánchez, A. J. Sáez-Castillo, and M. J. Olmo-Jiménez. A generalization of the betabinomial distribution. Applied Statistics, 56(1):51–61, January 2007.

Elliott:2007:UMM

[297] Michael R. Elliott and Nicolas Stettler.

Using a mixture model for multiple imputation in the presence of outliers: the ‘Healthy for life’ project. Applied Statistics, 56(1):63–78, January 2007.

Yuan:2007:MBE

[298] Ying Yuan and Roderick J. A. Little.

Model-based estimates of the finite population mean for two-stage cluster samples with unit non-response. Applied Statistics, 56(1):79–97, January 2007.

Fryzlewicz:2007:GXR

[299] Piotr Fryzlewicz, Véronique Delouille, and Guy P. Nason. GOES-8 X-ray sensor variance stabilization using the multiscale data-driven Haar–Fisz transform. Applied Statistics, 56(1):99–116, January 2007.

Spiess:2007:CET

[300] Martin Spiess. Corrigendum: Estimation of a two-equation panel model with mixed continuous and ordered categorical outcomes and missing data. Applied Statistics, 56(1):117, January 2007.

laCruz-Mesia:2007:SBC

[301] Rolando De la Cruz-Mesía, Fernando A. Quintana, and Peter Müller. Semiparametric Bayesian classification with longitudinal markers. Applied Statistics, 56 (2):119–137, March 2007. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ji:2007:BMM

[302] Yuan Ji, Guosheng Yin, Kam-Wah Tsui, Mikhail G. Kolonin, Jessica Sun, Wadih Arap, Renata Pasqualini, and Kim-Anh Do. Bayesian mixture models for complex high dimensional count data in phage display experiments. Applied Statistics, 56(2):139–152, March 2007.

Zhang:2007:RAR

[303] Lanju Zhang and William F. Rosenberger. Response-adaptive randomization for survival trials: the parametric approach. Applied Statistics, 56(2):153– 165, March 2007.

Karvanen:2007:EDB

[304] Juha Karvanen, Juha J. Vartiainen, Andrey Timofeev, and Jukka Pekola. Experimental designs for binary data in switching measurements on superconducting Josephson junctions. Applied Statistics, 56(2):167–181, March 2007.

Gryparis:2007:SLV

[305] Alexandros Gryparis, Brent A. Coull, Joel Schwartz, and Helen H. Suh. Semiparametric latent variable regression models for spatiotemporal modelling of mobile source particles in the greater Boston area. Applied Statistics, 56(2): 183–209, March 2007.

Baker:2007:EEP

[306] Stuart G. Baker, Constantine Frangakis, and Karen S. Lindeman. Estimating efficacy in a proposed randomized trial with initial and later non-compliance. Applied Statistics, 56(2):211–221, March 2007.

Wouters:2007:PDC

[307] Kristien Wouters, Abdellah Ahnaou, Jose Cortinas Abrahantes, Geert Molenberghs, Helena Geys, Luc Bijnens, and Wilhelmus H. I. M. Drinkenburg. Psychotropic drug classification based on sleep-wake behaviour of rats. Applied Statistics, 56(2):223–234, March 2007.

Ridall:2007:MUN

[308] P. G. Ridall, A. N. Pettitt, N. Friel, P. A. McCombe, and R. D. Henderson. Motor unit number estimation using reversible jump Markov chain Monte Carlo methods. Applied Statistics, 56(3): 235–269, May 2007. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Walls:2007:AEB

[309] Rebecca E. Walls, Stuart Barber, John T. Kent, and Mark S. Gilthorpe. An adaptive empirical Bayesian thresholding procedure for analysing microarray experiments with replication. Applied Statistics, 56(3):271–291, May

Paddock:2007:BVS

[310] Susan M. Paddock. Bayesian variable selection for longitudinal substance abuse treatment data subject to informative censoring. Applied Statistics, 56(3):293– 311, May 2007.

Smith:2007:CPR

[311] Brian J. Smith and Mary Kathryn Cowles. Correlating point-referenced radon and areal uranium data arising from a common spatial process. Applied Statistics, 56(3):313–326, May 2007. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Brezger:2007:AGM

[312] A. Brezger, L. Fahrmeir, and A. Hennerfeind. Adaptive Gaussian Markov random fields with applications in human brain mapping. Applied Statistics, 56(3): 327–345, May 2007.

Jones:2007:CSF

[313] Bradley Jones and Peter Goos. A candidate-set-free algorithm for generating D-optimal split-plot designs. Applied Statistics, 56(3):347–364, May 2007.

Bailey:2007:DTC

[314] R. A. Bailey. Designs for two-colour microarray experiments. Applied Statistics, 56(4):365–394, August 2007. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Butler:2007:TEE

[315] Adam Butler, Janet E. Heffernan, Jonathan A. Tawn, and Roger A. Flather. Trend estimation in extremes of synthetic North Sea surges. Applied Statistics, 56(4):395–414, August 2007.

Alshabani:2007:BAH

[316] A. K. S. Alshabani, I. L. Dryden, C. D.

Litton, and J. Richardson. Bayesian analysis of human movement curves. Applied Statistics, 56(4):415–428, August 2007.

Huggins:2007:NPE

[317] Richard Huggins and Wen-Han Hwang.

Non-parametric estimation of population size from capture-recapture data when the capture probability depends on a covariate. Applied Statistics, 56(4): 429–443, August 2007.

Arnold:2007:BFC

[318] Gillian M. Arnold, John C. Gower, Sugnet Gardner-Lubbe, and Niël J. Le Roux. Biplots of free-choice profile data in generalized orthogonal Procrustes analysis. Applied Statistics, 56 (4):445–458, August 2007.

Stoica:2007:TDO

[319] Radu S. Stoica, Vicent J. Martínez, and Enn Saar. A three-dimensional object point process for detection of cosmic filaments. Applied Statistics, 56(4):459– 477, August 2007.

Ricard:2007:SIO

[320] I. Ricard and A. C. Davison. Statistical inference for olfactometer data. Applied Statistics, 56(4):479–492, August 2007.

Shirley:2007:TST

[321] Eryl Shirley. Tests for a simple tree order restriction with application to doseresponse studies. Applied Statistics, 56 (4):493–494, August 2007.

Peddada:2007:ART

[322] Shyamal Peddada and Joe Haseman. Authors’ response: Tests for a simple tree order restriction with application to dose-response studies. Applied Statistics, 56(4):494–495, August 2007.

Nadarajah:2007:GBB

[323] Saralees Nadarajah. A generalization of the beta-binomial distribution. Applied Statistics, 56(4):496, August 2007.

Rodriguez-Avi:2007:ARG

[324] J. Rodríguez-Avi, A. Conde-Sánchez, A. J. Sáez-Castillo, and M. J. Olmo-Jiménez. Authors’ response: A generalization of the beta-binomial distribution. Applied Statistics, 56(4):496–497, August 2007.

Diggle:2007:ALD

[325] Peter Diggle, Daniel Farewell, and Robin Henderson. Analysis of longitudinal data with drop-out: objectives, assumptions and a proposal. Applied Statistics, 56(5):499–550, November 2007. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Forster:2007:BDR

[326] Jonathan J. Forster and Emily L. Webb. Bayesian disclosure risk assessment: predicting small frequencies in contingency tables. Applied Statistics, 56(5):551–570, November 2007.

Faraway:2007:MTD

[327] Julian J. Faraway, Matthew P. Reed, and Jing Wang. Modelling threedimensional trajectories by using Bézier curves with application to hand motion. Applied Statistics, 56(5):571–585, November 2007.

Hart:2007:APP

[328] Jeffrey D. Hart, Chris Koen, and Fred Lombard. An analysis of pulsation periods of long period variable stars. Applied Statistics, 56(5):587–606, November 2007.

Whitrow:2007:AOA

[329] Chris Whitrow. Algorithms for optimal allocation of bets on many simultaneous events. Applied Statistics, 56(5): 607–623, November 2007.

Anonymous:2007:IAV

[330] Anonymous. Index of authors, volume 56, 2007. Applied Statistics, 56(5):625– 627, November 2007.

Anonymous:2007:CV

[331] Anonymous. Contents of volume 56, 2007. Applied Statistics, 56(5):629–630, November 2007.

Chi:2008:ROC

[332] Yueh-Yun Chi and Xiao-Hua Zhou. Receiver operating characteristic surfaces in the presence of verification bias. Applied Statistics, 57(1):1–23, February 2008. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Heron:2008:CLS

[333] Elizabeth A. Heron and Cathal D.

Walsh. A continuous latent spatial model for crack initiation in bone cement. Applied Statistics, 57(1):25–42, February 2008.

Brentnall:2008:SMT

[334] Adam R. Brentnall, Martin J. Crowder, and David J. Hand. A statistical model for the temporal pattern of individual automated teller machine withdrawals. Applied Statistics, 57(1):43–59, February 2008.

Miglioretti:2008:MDM

[335] Diana L. Miglioretti and Elizabeth R. Brown. A marginalized diffusion model for estimating age at first lower endoscopy use from current status data. Applied Statistics, 57(1):61–74, February 2008.

Natarajan:2008:VEC

[336] Sundar Natarajan, Stuart R. Lipsitz, Garrett Fitzmaurice, Charity G. Moore, and Rene Gonin. Variance estimation in complex survey sampling for generalized linear models. Applied Statistics, 57(1): 75–87, February 2008.

Lloyd:2008:AML

[337] Chris J. Lloyd and Donald J. Frommer. An application of multinomial logistic regression to estimating performance of a multiple-screening test with incomplete verification. Applied Statistics, 57 (1):89–102, February 2008.

Thompson:2008:CIR

[338] J. R. Thompson, C. Minelli, K. R. Abrams, A. Thakkinstian, and J. Attia. Combining information from related meta-analyses of genetic association studies. Applied Statistics, 57(1): 103–115, February 2008.

Wichert:2008:SNP

[339] Laura Wichert and Ralf A. Wilke. Simple non-parametric estimators for unemployment duration analysis. Applied Statistics, 57(1):117–126, February 2008.

Reithinger:2008:MPP

[340] Florian Reithinger, Wolfgang Jank, Gerhard Tutz, and Galit Shmueli. Modelling price paths in on-line auctions: smoothing sparse and unevenly sampled curves by using semiparametric mixed models. Applied Statistics, 57(2):127–148, April 2008. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Schipper:2008:GMF

[341] Matthew Schipper, Jeremy M. G. Taylor, and Xihong Lin. Generalized monotonic functional mixed models with application to modelling normal tissue complications. Applied Statistics, 57(2): 149–163, April 2008.

Cappellari:2008:ELP

[342] Lorenzo Cappellari and Stephen P.

Jenkins. Estimating low pay transition probabilities accounting for endogenous selection mechanisms. Applied Statistics, 57(2):165–186, April 2008.

Rappold:2008:MMD

[343] Ana G. Rappold, Alan E. Gelfand, and David M. Holland. Modelling mercury deposition through latent spacetime processes. Applied Statistics, 57(2): 187–205, April 2008.

Hanson:2008:MSO

[344] Timothy E. Hanson, Athanasios Kottas, and Adam J. Branscum. Modelling stochastic order in the analysis of receiver operating characteristic data: Bayesian non-parametric approaches. Applied Statistics, 57(2):207– 225, April 2008.

Landajo:2008:MFP

[345] Manuel Landajo, Javier De Andrés, and Pedro Lorca. Measuring firm performance by using linear and nonparametric quantile regressions. Applied Statistics, 57(2):227–250, April 2008.

Paddock:2008:CBV

[346] Susan M. Paddock. Corrigendum: Bayesian variable selection for longitudinal substance abuse treatment data subject to informative censoring. Applied Statistics, 57(2):251–252, April 2008.

Soubeyrand:2008:MSS

[347] Samuel Soubeyrand, Leonhard Held, Michael Höhle, and Ivan Sache. Modelling the spread in space and time of an airborne plant disease. Applied Statistics, 57(3):253–272, June 2008. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Abayomi:2008:DMI

[348] Kobi Abayomi, Andrew Gelman, and Marc Levy. Diagnostics for multivariate imputations. Applied Statistics, 57 (3):273–291, June 2008.

Whitcher:2008:MRC

[349] Brandon Whitcher, Thomas C. M. Lee, Jeffrey B. Weiss, Timothy J. Hoar, and Douglas W. Nychka. A multi-resolution census algorithm for calculating vortex statistics in turbulent flows. Applied Statistics, 57(3):293–312, June 2008.

Palmer:2008:BSM

[350] Mark J. Palmer and Grant B. Douglas. A Bayesian statistical model for end member analysis of sediment geochemistry, incorporating spatial dependences.

Applied Statistics, 57(3):313–327, June 2008.

Bozza:2008:PEH

[351] Silvia Bozza, Franco Taroni, Raymond Marquis, and Matthieu Schmittbuhl. Probabilistic evaluation of handwriting evidence: likelihood ratio for authorship. Applied Statistics, 57(3):329–341, June 2008.

Glasbey:2008:SAR

[352] C. A. Glasbey and D. J. Allcroft. A spatiotemporal auto-regressive moving average model for solar radiation. Applied Statistics, 57(3):343–355, June 2008.

Nikolov:2008:SME

[353] Margaret C. Nikolov, Brent A. Coull, Paul J. Catalano, Edgar Diaz, and John J. Godleski. Statistical methods to evaluate health effects associated with major sources of air pollution: a casestudy of breathing patterns during exposure to concentrated Boston air particles. Applied Statistics, 57(3):357–378, June 2008.

Palmer:2008:MCG

[354] Karen J. Palmer, Martin S. Ridout, and Byron J. T. Morgan. Modelling cell generation times by using the tempered stable distribution. Applied Statistics, 57(4):379–397, September 2008. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Haslett:2008:SMP

[355] John Haslett and Andrew Parnell.

A simple monotone process with application to radiocarbon-dated depth chronologies. Applied Statistics, 57(4): 399–418, September 2008.

Quintana:2008:SBM

[356] Fernando A. Quintana, Peter Müller, Gary L. Rosner, and Mary V. Relling. A semiparametric Bayesian model for repeatedly repeated binary outcomes. Applied Statistics, 57(4):419–431, September 2008.

Morgan:2008:NMM

[357] B. J. T. Morgan and M. S. Ridout. A new mixture model for capture heterogeneity. Applied Statistics, 57(4):433– 446, September 2008.

Chen:2008:LPE

[358] Feng Chen, Richard M. Huggins, Paul S. F. Yip, and K. F. Lam. Local polynomial estimation of Poisson intensities in the presence of reporting delays. Applied Statistics, 57(4):447–459, September 2008.

Faddy:2008:EPP

[359] M. J. Faddy and D. M. Smith. Extended Poisson process modelling of dilution series data. Applied Statistics, 57(4):461– 471, September 2008.

Pauli:2008:AHW

[360] Francesco Pauli and Laura Rizzi. Analysis of heat wave effects on health by using generalized additive model and bootstrap-based model selection. Applied Statistics, 57(4):473–485, September 2008.

Hurn:2008:BHM

[361] Merrilee Hurn, Peter J. Green, and Fahimah Al-Awadhi. A Bayesian hierarchical model for photometric red shifts. Applied Statistics, 57(4):487–504, September 2008.

Butler:2008:LGM

[362] Adam Butler and Chris Glasbey. A latent Gaussian model for compositional data with zeros. Applied Statistics, 57(5):505–520, December 2008. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kaciroti:2008:BML

[363] Niko A. Kaciroti, Trivellore E. Raghunathan, M. Anthony Schork, and Noreen M. Clark. A Bayesian model for longitudinal count data with nonignorable dropout. Applied Statistics, 57(5):521–534, December 2008. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Clark:2008:DAC

[364] Robert G. Clark and Tanya C. Strevens.

Design and analysis of clustered, unmatched resource selection studies. Applied Statistics, 57(5):535–551, December 2008.

Sutradhar:2008:AIC

[365] Rinku Sutradhar and Richard J. Cook.

Analysis of interval-censored data from clustered multistate processes: application to joint damage in psoriatic arthritis. Applied Statistics, 57(5):553–566, December 2008.

Coetzer:2008:DRS

[366] R. L. J. Coetzer, R. F. Rossouw, and D. K. J. Lin. Dual response surface optimization with hard-to-control variables for sustainable gasifier performance. Applied Statistics, 57(5):567–587, December 2008.

Hu:2008:STS

[367] Zonghui Hu, Jing Qin, and Dean Follmann. Semiparametric two-sample changepoint model with application to human immunodeficiency virus studies. Applied Statistics, 57(5):589–607, December 2008.

King:2008:IDP

[368] Ruth King, Stephen P. Brooks, Chiara Mazzetta, Stephen N. Freeman, and Byron J. T. Morgan. Identifying and diagnosing population declines: a Bayesian assessment of lapwings in the UK. Applied Statistics, 57(5):609–632, December 2008.

Anonymous:2008:IAV

[369] Anonymous. Index of authors, volume 57, 2008. Applied Statistics, 57(5):633– 635, December 2008.

Anonymous:2008:CV

[370] Anonymous. Contents of volume 57, 2008. Applied Statistics, 57(5):637–638, December 2008.

Gilmour:2009:RE

[371] S. G. Gilmour and C. J. Skinner. Report of the Editors. Applied Statistics, 58(1): 1, February 2009. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Peng:2009:BHD

[372] Roger D. Peng, Francesca Dominici, and Leah J. Welty. A Bayesian hierarchical distributed lag model for estimating the time course of risk of hospitalization associated with particulate matter air pollution. Applied Statistics, 58(1): 3–24, February 2009.

Eastoe:2009:MNS

[373] Emma F. Eastoe and Jonathan A. Tawn.

Modelling non-stationary extremes with application to surface level ozone. Applied Statistics, 58(1):25–45, February 2009.

Holan:2009:HBM

[374] Scott H. Holan, Ginger M. Davis, Mark L. Wildhaber, Aaron J. DeLonay, and Diana M. Papoulias. Hierarchical Bayesian Markov switching models with application to predicting spawning success of shovelnose sturgeon. Applied Statistics, 58(1):47–64, February 2009.

Liu:2009:JAC

[375] Lei Liu and Xuelin Huang. Joint analysis of correlated repeated measures and recurrent events processes in the presence of death, with application to a study on acquired immune deficiency syndrome. Applied Statistics, 58(1):65–81, February 2009.

Imai:2009:SAR

[376] Kosuke Imai. Statistical analysis of randomized experiments with non-ignorable missing binary outcomes: an application to a voting experiment. Applied Statistics, 58(1):83–104, February 2009.

Hepworth:2009:DEP

[377] Graham Hepworth and Ray Watson.

Debiased estimation of proportions in group testing. Applied Statistics, 58(1): 105–121, February 2009.

Brabec:2009:SMN

[378] Marek Brabec, Ondřej Konár, Marek Malý, Emil Pelikán, and Jiří Vondráček.

A statistical model for natural gas standardized load profiles. Applied Statistics, 58(1):123–139, February 2009.

Butler:2009:CLG

[379] Adam Butler and Chris Glasbey. Corrigendum: A latent Gaussian model for compositional data with zeros. Applied Statistics, 58(1):141, February 2009.

Metcalf:2009:UBI

[380] C. Jessica E. Metcalf, David A. Stephens, Mark Rees, Svata M. Louda, and Kathleen H. Keeler. Using Bayesian inference to understand the allocation of resources between sexual and asexual reproduction. Applied Statistics, 58(2): 143–170, May 2009. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Beunckens:2009:MSA

[381] Caroline Beunckens, Cristina Sotto, Geert Molenberghs, and Geert Verbeke. A multifaceted sensitivity analysis of the Slovenian public opinion survey data. Applied Statistics, 58(2):171–196, May 2009.

Osborne:2009:EBM

[382] Jason A. Osborne and Tony E. Grift. Mestimation of Boolean models for particle flow experiments. Applied Statistics, 58(2):197–210, May 2009. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Yin:2009:BDF

[383] Guosheng Yin and Ying Yuan. Bayesian dose finding in oncology for drug combinations by copula regression. Applied Statistics, 58(2):211–224, May 2009.

Liu:2009:PNM

[384] Changying A. Liu and Thomas M.

Braun. Parametric non-mixture cure models for schedule finding of therapeutic agents. Applied Statistics, 58(2):225– 236, May 2009.

Lin:2009:UBV

[385] Yan Lin, Stuart Lipsitz, Debajyoti Sinha, Atul A. Gawande, Scott E. Regenbogen, and Caprice C. Greenberg. Using Bayesian p-values in a 2 × 2 table of matched pairs with incompletely classified data. Applied Statistics, 58(2): 237–246, May 2009.

Sisson:2009:TAM

[386] S. A. Sisson and Y. Fan. Towards automating model selection for a markrecapture-recovery analysis. Applied Statistics, 58(2):247–266, May 2009.

Gattone:2009:FAD

[387] Stefano A. Gattone and Tonio Di Battista. A functional approach to diversity profiles. Applied Statistics, 58(2): 267–284, May 2009.

Sangalli:2009:EET

[388] Laura M. Sangalli, Piercesare Secchi, Simone Vantini, and Alessandro Veneziani. Efficient estimation of threedimensional curves and their derivatives by free-knot regression splines, applied to the analysis of inner carotid artery centrelines. Applied Statistics, 58(3): 285–306, July 2009. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lehrach:2009:SBV

[389] Wolfgang P. Lehrach and Dirk Husmeier. Segmenting bacterial and viral DNA sequence alignments with a transdimensional phylogenetic factorial hidden Markov model. Applied Statistics, 58(3):307–327, July 2009.

Copas:2009:RPM

[390] John Copas and Claudia Lozada-Can. The radial plot in meta-analysis: approximations and applications. Applied Statistics, 58(3):329–344, July 2009.

Chagneau:2009:DPS

[391] P. Chagneau, F. Mortier, and N. Picard. Designing permanent sample plots by using a spatially hierarchical matrix population model. Applied Statistics, 58 (3):345–367, July 2009.

Bond:2009:LEL

[392] Simon J. Bond and Vernon T. Farewell. Likelihood estimation for a longitudinal negative binomial regression model with missing outcomes. Applied Statistics, 58 (3):369–382, July 2009.

Fouskakis:2009:PBR

[393] D. Fouskakis, I. Ntzoufras, and D. Draper. Population-based reversible jump Markov chain Monte Carlo methods for Bayesian variable selection and evaluation under cost limit restrictions. Applied Statistics, 58(3):383–403, July 2009.

Ailliot:2009:STM

[394] Pierre Ailliot, Craig Thompson, and Peter Thomson. Space-time modelling of precipitation by using a hidden Markov model and censored Gaussian distributions. Applied Statistics, 58(3):405–426, July 2009.

Koopman:2009:STC

[395] Siem Jan Koopman and Kai Ming Lee.

Seasonality with trend and cycle interactions in unobserved components models. Applied Statistics, 58(4):427–448, September 2009. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hout:2009:ELE

[396] Ardo Van Den Hout, Carol Jagger, and Fiona E. Matthews. Estimating life expectancy in health and ill health by using a hidden Markov model. Applied Statistics, 58(4):449–465, September 2009.

Gilmour:2009:ADN

[397] Steven G. Gilmour and Peter Goos.

Analysis of data from non-orthogonal multistratum designs in industrial experiments. Applied Statistics, 58(4):467– 484, September 2009.

Sanso:2009:SCC

[398] Bruno Sansó and Chris Forest. Statistical calibration of climate system properties. Applied Statistics, 58(4):485–503, September 2009.

Leiby:2009:IMR

[399] Benjamin E. Leiby, Mary D. Sammel, Thomas R. Ten Have, and Kevin G. Lynch. Identification of multivariate responders and non-responders by using Bayesian growth curve latent class models. Applied Statistics, 58(4):505–524, September 2009.

Guillot:2009:SPW

[400] Gilles Guillot, Niklas Lorén, and Mats Rudemo. Spatial prediction of weed intensities from exact count data and image-based estimates. Applied Statistics, 58(4):525–542, September 2009.

Spencer:2009:OMT

[401] Neil H. Spencer. Overcoming the multiple-testing problem when testing randomness. Applied Statistics, 58(4): 543–553, September 2009.

Najita:2009:NAB

[402] Julie S. Najita, Yi Li, and Paul J. Catalano. A novel application of a bivariate regression model for binary and continuous outcomes to studies of fetal toxicity. Applied Statistics, 58(4):555–573, September 2009.

Beunckens:2009:CMS

[403] Caroline Beunckens, Cristina Sotto, Geert Molenberghs, and Geert Verbeke. Corrigendum: A multifaceted sensitivity analysis of the Slovenian public opinion survey data. Applied Statistics, 58(4): 575–576, September 2009.

Binder:2009:SAA

[404] Harald Binder and Willi Sauerbrei. Stability analysis of an additive spline model for respiratory health data by using knot removal. Applied Statistics, 58(5):577–600, November 2009. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Keef:2009:SRA

[405] Caroline Keef, Jonathan Tawn, and Cecilia Svensson. Spatial risk assessment for extreme river flows. Applied Statistics, 58(5):601–618, November 2009.

Mair:2009:MWM

[406] Patrick Mair and Marcus Hudec. Multivariate Weibull mixtures with proportional hazard restrictions for dwell-time-based session clustering with incomplete data. Applied Statistics, 58(5):619–639, November 2009.

Robins:2009:RSI

[407] P. Robins, V. E. Rapley, and N. Green.

Realtime sequential inference of static parameters with expensive likelihood calculations. Applied Statistics, 58(5): 641–662, November 2009.

Li:2009:ETH

[408] X. Li, B. N. Thomas, S. M. Rich, D. Ecker, J. K. Tumwine, and A. S. Foulkes. Estimating and testing haplotype-trait associations in nondiploid populations. Applied Statistics, 58(5):663–678, November 2009.

Tomassini:2009:SAE

[409] Lorenzo Tomassini, Peter Reichert, Hans R. Künsch, Christoph Buser, Reto Knutti, and Mark E. Borsuk. A smoothing algorithm for estimating stochastic, continuous time model parameters and its application to a simple climate model. Applied Statistics, 58(5):679– 704, November 2009.

Jones:2009:ECB

[410] Geoffrey Jones, Wesley O. Johnson, and W. Daan Vink. Evaluating a continuous biomarker for infection by using observed disease status with covariate effects on disease. Applied Statistics, 58 (5):705–717, November 2009.

Yuan:2009:BDF

[411] Ying Yuan and Guosheng Yin. Bayesian dose finding by jointly modelling toxicity and efficacy as time-to-event outcomes. Applied Statistics, 58(5):719– 736, November 2009.

Bowman:2009:SSS

[412] Adrian W. Bowman, Marco Giannitrapani, and E. Marian Scott. Spatiotemporal smoothing and sulphur dioxide trends over Europe. Applied Statistics, 58(5):737–752, November 2009.

Anonymous:2009:IAV

[413] Anonymous. Index of authors, volume 58, 2009. Applied Statistics, 58(5):753– 755, November 2009.

Anonymous:2009:CV

[414] Anonymous. Contents of volume 58, 2009. Applied Statistics, 58(5):757–758, November 2009.